### v1

# pMixFed: Efficient Personalized Federated Learning through Adaptive Layer-Wise Mixup

Abstract
Traditional Federated Learning (FL) methods encounter significant challenges when dealing with heterogeneous data and providing personalized solutions for non-IID scenarios. Personalized Federated Learning (PFL) approaches aim to address these issues by balancing generalization and personalization, often through parameter decoupling or partial models that freeze some neural network layers for personalization while aggregating other layers globally. However, existing methods still face challenges of global-local model discrepancy, client drift, and catastrophic forgetting, which degrade model accuracy. To overcome these limitations, we propose pMixFed, a dynamic, layer-wise PFL approach that integrates mixup between shared global and personalized local models. Our method introduces an adaptive strategy for partitioning between personalized and shared layers, a gradual transition of personalization degree to enhance local client adaptation, improved generalization across clients, and a novel aggregation mechanism to mitigate catastrophic forgetting. Extensive experiments demonstrate that pMixFed outperforms state-of-the-art PFL methods, showing faster model training, increased robustness, and improved handling of data heterogeneity under different heterogeneous settings. Our code is available for reproducing our results: https://github.com/YasMinSdt/pMixFed.
1 Introduction
Figure 1: Discrepancy between personalized and global shared layers in Partial PFL: (1) The global model, $G^{t}$, is constructed by aggregating asynchronous local updates from clients, denoted as $L^{t}_{i}$, $L^{t}_{j}$, and $L^{t}_{k}$. (2),(3) In communication round $t$, available clients $i$ and $j$ aggregate shared parameters to produce the updated global model $G^{t+1}$, while the personalized parameters, such as $L^{t}_{k}$, remain unchanged for unavailable clients. (4) This integration of distinct models, $G^{t+1}$ and $L^{t}_{k}$, induces inconsistencies in the overall model updates. (Bottom) During the joint training of generalized and personalized models, the gradient updates from the generalized layers are impacted by the gradients from personalized layers, resulting in catastrophic forgetting, performance drop and slower convergence rates.
The goal in federated learning (FL) [22] is to facilitate collaborative learning of several machine learning (ML) models in a decentralized scheme. FL requires addressing data privacy, catastrophic forgetting, and client drift problem  [^1] A phenomenon where the global model fails to serve as an accurate representation because local models gradually drift apart due to high data heterogeneity. [38, 15, 43, 31, 36]. Existing FL methods cannot address all these challenges with non-IID (non-Independent and Identically Distributed) data. For instance, although “FedAvg”  [32] demonstrates strong generalization performance, it fails to provide personalized solutions for a cohort of clients with non-IID datasets. Hence, the global model, or one “average client” in “FedAvg”, may not adequately represent all individual local models in non-IID settings due to client-drift [53]. Personalized FL (PFL) methods handle data heterogeneity by considering both generalization and personalization during the training stage. Since there is a trade-off between generalization and personalization in heterogeneous environments, PFL methods leverage heterogeneity and diversity as advantages rather than adversities [35, 47]. A group of PFL approaches train personalized local models on each device while collaborating toward a shared global model. Partial PFL, also known as parameter decoupling, involves using a partial model sharing, where only a subset of the model is shared while other parameters remain “frozen” to balance generalization and personalization until the subsequent round of local training.
While partial PFL methods are effective in mitigating catastrophic forgetting, strengthening privacy, and reducing computation and communication overhead [34, 45], there are still some unaddressed challenges. First, the question of when, where, and how to optimally partition the full model? remains unresolved. Recent studies [34, 45] have shown that there is no “one-size-fits-all” solution; the best or optimal partitioning strategy depends on factors such as task type (e.g., next word prediction or speech recognition) and local model architecture. An improper partitioning choice can lead to issues such as underfitting, overfitting, increased bias, and catastrophic forgetting. Some studies [29] suggest that personalized layers should reside in the base layers, while others [9, 4] argue that the base layers contain more generalized information and should be shared. Further, the use of a fixed partitioning strategy across all communication rounds for heterogeneous clients can limit the efficacy of collaborative learning. For instance, if the performance of the client suddenly drops due to new incoming data, the partitioning strategy should be changed because the client requires more frozen layers. Another issue is catastrophic forgetting of the previously shared global knowledge after only a few rounds of local training because the shared global model can be completely overwritten by local updates leading to generalization degradation [31, 41, 15, 54]. Most importantly, partial models may experience slower convergence compared to full model personalization, as frozen local model updates can diverge in an opposite direction from the globally-shared model. Since the generalized and personalized models are trained on non-IID datasets, there might also be a domain shift, leading to model discrepancy as depicted in Figure 1. These discrepancies arise from variations in local and global objective functions, differences in initialization, and asynchronous updates [56, 24]. As a result, merging the shared and the personalized layers can disrupt information flow within the network, impede the learning process, and lead to a slower convergence rate or accuracy drop in partial PFL models such as FedAlt and FedSim [34] [^2] More details on this is discussed in section 4.. Further, while partial PFL techniques contribute to an overall improved training accuracy, they can reduce the test accuracy on some devices, particularly in devices with limited samples, leading to variations in results in terms of the performance level [34]. Hence, there is a need for novel solutions to achieve the following goals in PFL:
  • Dynamic and Adaptive Partitioning: The balance between shared and personalized layers should be dynamically and adaptively adjusted for each client during every communication round, rather than relying on a static, fixed partitioning strategy for all participants.
  • Gradual Personalization Transition: The degree of personalization should transition gradually across layers, as opposed to an ”all-or-nothing” approach that employs strict partitioning or hard splits within the model discussed in Figure 1. This ability allows nuanced adaptation for individual client needs due to heterogeneity.
  • Improved Generalization Across All Clients: The average personalization accuracy should be such that the global model is unbiased toward specific subsets of clients.
  • Mitigation of Catastrophic Forgetting: The strategy should address the catastrophic forgetting problem by incorporating mechanisms to strengthen the generalization and retain the state of the previous global model when updating the global model in aggregation.
  • Scalability and Adaptability: The approach should be fast, scalable, and easily adaptable to new cold-start clients while accounting for model/device heterogeneity.
To achieve the above, we propose “pMixFed”, a layer-wise, dynamic PFL approach that integrates Mixup [61] between the shared global and personalized local models’ layers during both the broadcasting (global model sharing with local clients) and aggregation (aggregating distributed local models to update the global model) stages within a partial PFL framework. Our main contributions include:
  • We develop an online, dynamic interpolation method between local and global models using Mixup [57], effectively addressing data heterogeneity and scalable across varying cohort sizes, degrees of data heterogeneity, and diverse model sizes and architectures.
  • Our solutions facilitate a gradual increase in the degree of personalization across layers, rather than relying on a strict cut-off layer, helping to mitigate client drift.
  • We introduce a new fast and efficient aggregation technique which addresses catastrophic forgetting by keeping the previous global model state.
  • “pMixFed” reduces the participant gap (test accuracy for cold-start users) and the out-of-sample gap (test accuracy on unseen data) caused by data heterogeneity through linear interpolation between client updates, thereby mitigating the impact of client drift.
2 Related Work
PFL aims to adapt local models to the individual needs, preferences, and contexts of each participant. Since the inception of FL, various PFL approaches have been explored to address the challenges stemming from heterogeneity, categorized into three areas: 1) data-centric strategies 2) global model adaptation, and 3) local model personalization. The following provide a more detailed discussion of these three approaches:
Data-centric: These approaches in PFL aim to address data heterogeneity and class imbalances by manipulating statistical data properties, such as data size, distribution, and local data selection [47]. Examples of these techniques include data normalization, feature engineering, data augmentation, synthetic data generation, and client selection. Representative methods, such as Astraea [11], P2P k-SMOTE [50], FedMCCS [1], FedAug [19], [63], and FedHome [52], exemplify data-centric approaches. The shortcoming of these methods is that they often modify the natural distribution and statistical characteristics of federated data, potentially injecting bias or eliminating valuable, rare information. Additionally, many of these strategies depend on the availability of proxy data on the global server, which may conflict with the privacy regulations of some entities.
Global Model Adaptation: In these approaches, a single global model is maintained on the server and subsequently adapted to individual local models in the following phase. The primary goals of these methods are: 1) learning a robust, generalized model and 2) enabling fast, efficient local model adaptation. Several techniques employ regularization terms to mitigate client drift and prevent model divergence during local updates e.g., FedProx [27], SCAFFOLD [21], FL-MOON [26], and FedCurv [42]. Some of these approaches such as PerFedAvg [12], pFedMe [46], and pFedHN [39] leverage Meta-learning, and transfer learning e.g., FedStag [55], and DPFed [58], yet they introduce specific challenges. Meta-learning methods can be computationally intensive, while regularization techniques may add computational overhead by incorporating additional terms in the objective function. Similarly, transfer learning algorithms can be inefficient in terms of communication overhead and often require a public dataset to enhance the global model on the server. A common limitation across most adaptation techniques is the need for uniform model architecture across all clients, requiring devices with varying computational capabilities to use the same model size.
Local Model Personalization: The limitations of existing PFL methods have led to the development of local model personalization approaches that focus on training customized models for each client. Some approaches utilize multi-task learning (MTL), a collaborative framework that facilitates information exchange across distinct tasks (e.g., FedRes [2], MOCHA [44], FedAMP [16], and Ditto [28]). Another category leverages knowledge distillation (KD) to support personalization when client-specific training objectives differ, as seen in FedDF [30], FedMD [25], FedGen [64], and FedGKT [14]. However, both MTL and KD can incur significant computational and communication overhead, limiting their scalability for large-scale FL deployments across resource-constrained devices.
Partial PFL: These methods, also known as parameter decoupling, are the primary focus of this paper. These methods mitigate catastrophic forgetting by retaining shared components while freezing other layers for personalization. FedPer [4] introduced partial models in FL, sharing only initial layers with generalized information and reserving final layers for personalization. FedBABU [33] divides the network into a shared body and a frozen head with fully connected layers. Other frameworks, such as FURL [6] and LG-FedAvg [29], apply partial PFL by retaining private feature embeddings or using compact representation learning for high-level features, respectively. Two baseline methods in this paper are FedAlt [43] and FedSim [34]. FedAlt uses a stateless FL paradigm, reconstructing local models from the global model, while FedSim synchronously updates shared and local models with each iteration.
Limitations: These methods face challenges such as model update discrepancies (as shown in Figure 1) and catastrophic forgetting, where shared layers may undergo significant changes after only a few rounds of local training, resulting in sudden accuracy drops. Additionally, users with high personalization accuracy may freeze more layers than cold-start users or unreliable participants, who should rely more heavily on the global model. These challenges have motivated our development of a dynamic, adaptive layer-wise approach to balance generalization and personalization across clients, allowing tuning at different communication rounds to accommodate varying performance conditions.
3 Problem Formulation
Consider $K$ collaborating clients (or agents), each trying to optimize a local loss function $F_{k}(\theta)$ on the distributed local dataset ${D_{k}}={(x_{k},y_{k})}$, where $(x,y)$ shows the data features and the corresponding labels, respectively. Since the agents collaborate, the parameters $\theta$ (parameters of the global model) are shared across the agents. A basic FL objective function aims to optimize the overall global loss:
$\min_{\theta}F(\theta)=\sum_{k=1}^{K}\frac{|\mathcal{D}_{k}|}{|\mathcal{D}|}F_{k}(\theta)\quad\quad,$ (1)
$F_{k}(\theta)=\frac{1}{|\mathcal{D}_{k}|}\sum_{(x,y_{i})\in\mathcal{D}_{k}}\ell(f_{k}(\theta,x_{i}),y_{i})$
where $|\mathcal{D}|=\sum_{k=1}^{K}|\mathcal{D}_{k}|$ and $F(\theta)$ is the global loss function of FedAvg [32]. FL is performed in at iterative fashion. At each round, each client downloads the current version of the global model and trains it using their local data. Clients then send the updated model parameters to the central server. The central server uses the model updates from the selected clients and aggregates them to update the global model. Iterations continue until convergence.
In Equation 1, the assumption is that the data is collected from an IID distribution, and all clients should train their data according to the exact same model. However, this assumption cannot be applied within many practical FL settings due to the non-IID nature of data and resource limitations [18, 17]. In the PFL settings with high heterogeneity and non-IID data distribution, the same issue persists and the local parameters need to be customized toward each agent. PFL extends FL by solving the following [26, 34]:
$\min_{\theta,\theta_{k}}\sum_{k=1}^{K}\frac{1}{|\mathcal{D}_{k}|}(\mathcal{F}_{k}(\theta_{k})+\alpha_{k}\|\mathbf{\theta_{k}-\theta}\|^{2}).$ (2)
PFL explicitly handles data heterogeneity through the term $\mathcal{F}_{k}(\theta_{k})$ which accounts for model heterogeneity by considering personalized parameter $\theta_{k}$ for client $k$. Meanwhile, $\theta$ represents the shared global model parameters in Equation 2, and $\alpha_{k}$ acts as a regularizer and indicates the degree of personalization tuning collaborative learning between personalized local models $\theta_{k}$ and generalized global model $\theta$. When $\alpha_{k}$ is small, the personalization power of the local models will be increased. If $\alpha_{k}$ is large, the local models’ parameters tend to be closer to the global model parameters $\theta$.
3.1 Partial Personalized Model
The limitations of full model personalization methods with global and fully independent local models are discussed in Section 2. Partial PFL methods improve personalization by providing more flexibility through allowing clients to choose which parts of their models to be personalized based on their specific needs and constraints for improved performance. Let $L_{k}^{t}$ be a partial local model $k$ in round $t$ which is partitioned into two parts $\langle L_{l,k}^{t};L_{g,k}^{t}\rangle$, where $l,g\subseteq\{1,\dots M\}$ are the personalized and global layers, and $M$ is the number of layers. We can integrate both personalized and generalized layers in a local model $L_{k}$ as:
$\mathcal{F}_{k}(\theta_{k})=\ell(f_{k}(\langle L_{g,k};L_{l,k}\rangle,x_{k}),y_{k})$ (3)
Among different partitioning strategies for partial PFL [34], the most popular technique is to assign local personalized layers $L_{l,k}^{t}$ to final layers and allow the base layers $L_{g,k}^{t}$ to share the knowledge similar to FedPer [4]. This choice aligns with insights from MAML  [^3] Model-Agnostic Meta-Learning algorithm, suggesting that initial layers keeps general and broad information while personalized characteristics manifest prominently in the higher layers. Accordingly, we would have:
$\mathcal{F}_{k}(\theta_{k})=L_{l}^{(t)}(L_{g}^{(t)}(x_{k}))\xrightarrow{localupdate}L^{\prime}_{l}(L^{\prime}_{g}(x_{k}))\quad,$
$L^{\prime}_{l}(L^{\prime}_{g}(x_{k}))\xrightarrow{broadcasting}L_{l}^{(t+1)}(G^{(t+1)}(x_{k}))$
For simplicity $L_{g}=L_{g,k}$ and $L_{l}=L_{g,k}$ where $\{1\leq g\leq s\leq l\leq M\}$ and $s$ is the split(cut) layer. The objective in solving Equation 3.1 is to find the optimal $s$ (cut layer) which minimizes the personalization objective: $\sum_{k=1}^{K}\frac{1}{|\mathcal{D}_{k}|}\mathcal{F}_{k}(\theta_{k})=L_{l}^{(t)}(L_{g}^{(t)}(x_{k}))$.
In partial models, after several rounds of local training, both personalized and global layers of local model are updated. This update could be synchronous like FedSim or asynchronous as in FedAlt. The Personalized layers will be frozen until the next communication round, $L_{l}^{(t+1)}=L^{\prime}_{l}$ and the global layers will be sent to the server for global model aggregation : $G^{(t+1)}\leftarrow\sum_{k=1}^{K}\frac{|\mathcal{D}_{k}|}{|\mathcal{D}|}L^{\prime}_{g}(x_{k})$. In the next broadcasting phase, the shared layers of the local model will be updated as $L_{g}^{(t+1)}\leftarrow G^{(t+1)}$. Our goal is to benefit from Mixup to improve personalization.
4 The Proposed Method
4.1 Mixup
Mixup is a data augmentation technique for enhancing model generalization [61] based on learning to generalize on linear combinations of training examples. Variations of Mixup have consistently excelled in vision tasks, contributing to improved robustness, generalization, and adversarial privacy. Mixup creates augmented samples as:
$\small\begin{split}\bar{x}=\lambda.x_{i}+(1-\lambda).x_{j}\\ \bar{y}=\lambda.y_{i}+(1-\lambda).y_{j},\end{split}$ (4)
where $x_{i}$ and $x_{j}$ are two input samples, $y_{i}$ and $y_{j}$ are the corresponding labels, $\lambda$, $\lambda\sim\text{Beta}(\alpha,\alpha),\lambda\in[0,1]$, is the degree of interpolation between the two samples. Mixup relates data points belonging to different classes which has been shown to be successful in mitigating overfitting and improving model generalization [49, 13, 62]. There are many variants of Mixup that has been developed to address specific challenges and enhance its effectiveness. For instance, AlignMixup [48] improves local spatial alignment by introducing transformations that better preserve semantic consistency between input pairs. Manifold Mixup [49] extends the concept to hidden layer representations, acting as a powerful regularization technique by training deep neural networks (DNNs) on linear combinations of intermediate features. CutMix [60] replaces patches between images, blending visual information while retaining spatial structure. Remix [7] addresses class imbalance by assigning higher weights to minority classes during the mixing process, enhancing the robustness of the trained model on imbalanced datasets. AdaMix [13] dynamically optimizes the mixing distributions, reducing overlaps and improving training efficiency. This linear interpolation also serves as a regularization technique that shapes smoother decision boundaries, thereby enhancing the ability of a trained model to generalize to unseen data. Mixup can also increase robustness against adversarial attacks [62, 5]; and improves performance against noise, corrupted labels, and uncertainty as it relaxes the dependency on specific information [13].
Figure 2: Workflow of pMixFed: Mixup is used in two stages. 1-Broadcasting: when transferring knowledge to local models, the frozen personalized model $L_{k}^{(t)}$ is mixed up with global model $G^{(t)}$ according to the adaptive mix factor $\mu_{k}^{(t)}$ which determines layer-wise mixup degree $\lambda_{i}$ for layer $i$. 2-Aggregation: The updated global model $G^{(t+1)}$ is generated through a Mixup between the updated local model $L^{(t+1)}$ and the current shared global model $G^{\prime(t)}$. Consequently, a state of the previous global model $G^{\prime(t)}$ is retained, which aids in mitigating catastrophic forgetting.
4.2 pMixFed : Partial Mixed up Personalized federated learning
Our goal is to leverage the well-established benefits of Mixup in the context of personalized FL. While Mixup has previously been employed in FL frameworks, such as XORMixup [40], FEDMIX [57], and FedMix [51], prior studies have primarily focused on using Mixup for data augmentation or data averaging. We propose pMixFed by integrating Mixup on the model parameter space, rather using it on the feature space. We apply Mixup between the parameters of the global and the local models in a layer-wise manner for more customized and adaptive PFL. Our approach eliminates the need for static and rigid partitioning strategies. Specifically, during both the broadcasting and aggregation stages of our partial PFL framework, we generate mixed model weights using a an interpolation strategy which is illustrated in Figure 2. Mixup offers the flexibility in combining models by introducing a mix degree for each layer $\lambda_{i}$, which changes gradually according to $\mu$, i.e., the mix factor. Parameter $\mu$ is also updated adaptively in each communication round and for each client according to the test accuracy of the global and the local model during the evaluation phase of FL. $\mu$ is computed as follows:
$\small\mu_{k}^{t}=1-\frac{1}{1+e^{-\delta(Acc^{b}-1)}}$ (5)
$\delta=(t/T)$, where $t$, and $T$ are the current communication round, and the overall number of communication rounds respectively. Also, $Acc$ calculated as: $Acc=(Acc_{k}^{t}/Acc_{overall}^{(t-1)})$ in the broadcasting phase and average test accuracy of the previous global model $\left(Acc=G^{t}(x,\theta^{t}),y\right)$ on all local test sets $x=\{x_{1},x_{2},\ldots,x_{K}\}$. , in the aggregation stage. $b$ is the offset parameter for sigmoid function which is set to $2$ after several experiments. More detailed discussion on parameter $\mu$’s rule of update is discussed in section 5 and the reason behind this formulation is discussed in the Appendix 8.
As shown in Figure 2, Mixup is applied in two distinct stages of FL. Firstly, when transferring shared knowledge to local models, the local model $L_{k}$ is mixed up with the current global model $G$ according to the dynamic mixing factor $\mu$, which determines the change ratio of $\lambda_{i}$ (layer-wise Mixup degree in Eq. (4)) across different layers. $\lambda_{i}$ gradually is changed from $1\rightarrow 0$ as we move from the head to the base layer. $\lambda_{i}=1$ means sharing the 100% of the global model and $\lambda_{i}=0$ means that the corresponding layer in local model is frozen and will not be mixed up with the global model $G$. Calculation of the Mixup degree of layer $i$ $\lambda_{i}$ at both broadcasting and aggregation stages is performed as follows:
$\small\begin{split}\text{Broadcasting Stage:}\quad\quad\lambda_{i}=\begin{cases}1&\lambda_{i}>1\\ \mu*(n-i)&\lambda_{i}\leq 1\end{cases}\\ \text{Aggregation Stage:}\quad\quad\lambda_{i}=\begin{cases}0&\lambda_{i}\leq 0\\ 1-(i*\mu)&\lambda_{i}>0,\end{cases}\end{split}$ (6)
where $n$ is total number of layers, $i$ is the current layer number starting from the base layer to the head as $i=0\rightarrow i=n$. $\mu$ is the mix factor which will be adaptively updated in each communication round $t$ and for each local model $k$ according to Eq. 5.
Algorithm 1 pMixFed: Broadcasting global to local model
1:Input: Initial states global model: $GM^{(0)}$, local models: $\{LM^{(0)}_{i}\}_{i=1,\ldots,M}$, Number of communication rounds $T$, number of devices per round $m$, Number of layers in local models $\{L_{i}\}$
2:for $t=0,1,\ldots,T-1$ do
3:    Server selects $K$ devices $S(t)\subset\{1,\ldots,N\}$
4:    Update $\mu$ for each $LM_{i}$, $i=\{1,\ldots,K\}$
5:    Server broadcasts $GM^{(t)}$ to each device in $S(t)$
6:    for each device $m\in S(t)$ in parallel do
7:        $(LM^{(t+1)}_{m},GM^{(t+1)}_{m})=\text{Mixup}[(LM^{\prime(t+1)}_{m},GM^{\prime(t+1)}_{m}),\mu]$
8:        Device sends $GM^{(t+1)}_{m}$ back to server
9:        Update $\mu_{i=1,\ldots,K}$
10:    end for
11:    Server updates $GM^{(t+1)}=\frac{1}{K}\sum_{i\in S(t)}GM^{(t+1)}_{m}$
12:end for
Algorithm 2 Proposed Mixup for Aggregation
1:Input: Initial states global model: $GM^{(0)}$, Number of communication rounds $T$, Number of local iterations $Itr$, number of devices per round $m$, Mixup degree $\lambda$
2:for $t=0,1,\ldots,T-1$ do
3:    Server broadcasts $GM^{(t)}$ to each device in $S(t)$: $LM_{i}^{(t)}$
4:    Update $\lambda_{i=1,\ldots,M}$ for each device
5:    for each device $i\in S(t)$ do
6:        for $epoch=0,1,\ldots,Itr-1$ do
7:           Train local model:
8:                         $LM_{i}^{(t+1)}=GM^{(t)}$
9:                     $GM^{(t+1)}=$
10:$\hskip 65.44133pt\text{Mixup}(LM_{i}^{(t+1)},GM^{(t)})$
11:        end for
12:        Device sends $GM^{(t+1)}$ back to server
13:        Adaptively update $\lambda_{i=1,\ldots,M}$
14:    end for
15:end for
4.2.1 Broadcasting : Global to local model transfer
This stage involves sharing global knowledge with local clients. In the existing PFL methods, the same weight allocation is typically applied to each heterogeneous local model. In our work, we personalize this process by allowing the local model to select the proportion of layers it requires. For instance, for a cold-start user, more information should be extracted from the shared knowledge model, implying that a few layers should be frozen for personalization. Accordingly, A history of the previous global model $G^{t}$ will remain which helps with catastrophic forgetting of the generalized model. Additionally, we introduce a gradual update procedure where the value of $\lambda$ gradually decreases from one (indicating fully shared layers) from the base layer to the end, based on the mixing factor $\mu$. The mix layer is adaptively updated in each communication round for each client individually, according to personalization accuracy. With this adaptive and flexible approach, not only can upcoming streaming unseen data be managed, but also the participation gap (test accuracy of new cold start users) would be improved. The Broadcasting phase is illustrated in 1 which is inspired by [20]. The update rule of local model $L_{k}^{(t)}$ is as follows:
$\small\begin{split}&{L^{\prime}}^{(t)}_{k,i}=(\lambda^{(t)}_{k,i}).G^{(t)}_{i}+(1-\lambda^{(t)}_{k,i}).L^{(t)}_{k,i}\\ &L_{k}^{(t+1)}={L^{\prime}}^{(t)}_{k}-\eta\nabla F_{k}({L^{\prime}}^{(t)}_{k}).\end{split}$ (7)
4.2.2 Aggregation: Local to Global Model Transfer
Existing methods primarily categorize layers into two types: personalized layers and generalized layers. The updated global layers from different clients, $G_{k}^{(t+1)}$, are typically aggregated using Equation 3, which can lead to catastrophic forgetting. Since The base layers of the global model serve as the backbone of shared knowledge [37], This issue arises because, during each local update, the generalized layers undergo substantial modifications [34, 33]. When the global model is updated by simply aggregating local models, valuable information from previously shared knowledge may be lost, leading to forgetting—even if $G^{t}$ performs better than the newly aggregated model. To address this challenge, we propose a new strategy that applies Mixup between the gradients of the previous global model and each client’s local model before aggregation. For each client $i$, the mixup coefficient $\lambda_{i}$ gradually increases from 0 to 1, moving from the head to the base layer, controlled by the mix factor $\mu_{i}$. Additionally, the base layer is adaptively updated based on the communication round and the generalization accuracy, ensuring robust integration of shared and personalized knowledge. It should be noted that parameter $\mu$ is constant in the aggregation stage for all clients as it’s dependent to the average performance of the previous global model.
$\small\begin{split}&{G^{\prime}}^{(t)}_{k,i}=(\lambda^{(t)}_{i,k}).G^{(t)}_{i}+(1-\lambda^{(t)}_{k,i}).L^{(t+1)}_{k,i},\\ &G^{(t+1)}=\sum_{k=1}^{K}\frac{|D_{k}|}{\sum_{k=1}^{K}|D_{k}|}G^{\prime(t+1)}_{k}.\end{split}$ (8)
The high-level block-diagram visualization of the proposed method is shown in Figure 2. It is important to note that the sizes of local models $LM^{(0)}_{i}$ can differ from each other. Consequently, the size of $GM^{(0)}$ should be greater than the maximum size of local models. The parameter $\lambda$ in Equation 4 determines the Mixup degree between the shared model $GM$ and the local models $LM^{(\cdot)}_{i}$, while $\mu$ governs the slope of the change in $\lambda$ across different layers. The degree of Mixup gradually decreases according to the parameter $\mu$ from 00 to $1$. In this scenario, $\lambda=0$ for the first base layer, indicating total sharing, while $\lambda=1$ applies to the final layer, which represents no sharing. The underlying concept is that the base layer contains more general information, whereas the final layers retain client-specific information. The use of the parameter $\mu$, relative to the number of local layers, eliminates the need for a specified cut layer $k$ and allows its application across different model sizes and layers. The parameter $\mu_{i}$ is adaptively updated based on the personalized and global model accuracy for each client. Algorithm 1 shows how Mixup is used as a shared aggregation technique between individual clients and the server. In each training round, only one client is Mixed up with the global model, and $\lambda$ is adaptively learned based on the objective function using online learning. Algorithm 2 shows how Mixup is employed as a shared aggregation technique between the clients and the server. In each training round, only one client is “Mixed up” with the global model and the $\lambda$ parameter is adaptively learned based on the objective function using online learning. For a theoretical convergence analysis of pMixFed, please refer to the Appendix 9.
5 Experiments
Table 1: Comparison of the State-of-The-Art Methods using Mobile-Net model
Method CIFAR-10 CIFAR-100
$N=100$ $N=10$ $N=100$ $N=10$
$C=10\%$ $C=100\%$ $C=10\%$ $C=100\%$ $C=10\%$ $C=100\%$ $C=10\%$ $C=100\%$
FedAvg 26.93 25.61 15.64 19.55 4.50 4.54 17.65 21.24
FedAlt 39.36 46.06 39.60 42.79 27.85 19.07 48.30 14.31
FedSim 51.12 45.43 40.30 35.43 28.18 19.52 47.41 48.30
FedBaBU 28.70 25.98 17.58 20.74 4.63 4.72 17.46 17.57
Ditto 26.58 62.77 18.60 43.98 7.00 16.07 16.89 20.74
Per-FedAvg 34.30 43.59 19.52 38.29 24.04 30.65 16.27 37.20
Lg-FedAvg 34.52 34.17 48.88 58.51 5.65 5.73 31.89 35.78
pMixFed 69.94 72.42 54.90 74.62 45.62 56.63 54.71 58.25
Table 2: Comparison of the State-of-The-Art Methods using CNN model
Method CIFAR-10 CIFAR-100 MNIST
$N=100$ $N=10$ $N=100$ $N=10$ $N=100$ $N=10$
$C=10\%$ $C=100\%$ $C=10\%$ $C=100\%$ $C=10\%$ $C=100\%$ $C=10\%$ $C=100\%$ $C=100\%$ $C=100\%$
FedAvg 54.78 56.82 44.11 54.37 25.77 26.73 34.47 39.93 97.54 98.59
FedAlt 56.41 56.77 69.50 64.80 15.19 10.56 28.30 26.53 97.37 99.21
FedSim 59.90 56.07 63.46 38.34 14.80 10.46 27.00 26.47 98.63 99.60
FedBaBU 53.12 54.60 39.77 53.21 16.77 17.33 25.60 32.47 98.19 99.07
Ditto 46.86 79.60 31.65 60.75 27.16 42.93 25.38 55.27 98.03 95.51
Per-FedAvg 39.37 45.03 10.00 48.13 32.67 39.01 8.71 41.21 98.32 50.34
Lg-FedAvg 62.28 62.99 62.46 71.73 28.75 28.03 33.75 45.76 97.65 98.82
pMixFed 65.30 75.49 74.36 75.06 34.66 41.56 43.47 51.46 99.88 99.98
5.1 Experimental Setup
Datasets: We used three datasets widely used federated learning: MNIST [23], CIFAR-10, and CIFAR-100 [3], We followed the setup in [32] and [33] to simulate heterogeneous, non-IID data distributions across clients for both train and test datasets. The maximum number of classes per user is set to $S=5$ for CIFAR-10 and MNIST, and $S=50$ for CIFAR-100. Experiments were conducted across varying heterogeneous settings, including small-scale ($N=10$) and large-scale ($N=100$) client populations, with different client participation rates $C=[100\%,10\%]$ to measure the effects of stragglers.
Baselines and backbone: We compare two version of our method (i) pMixFed: an adaptive and dynamic mixup-based PFL approach, and (ii) pMixFed-Dynamic: a dynamic-only mixup variant where the parameter $\mu$ is fixed across communication rounds, against several baselines. These baselines include: FedAvg [32], FedAlt [34], FedSim [34], FedBABU [33]. Additionally, we compare against full model personalization methods, Ditto [28], Per-FedAvg [12] , and LG-FedAvg [29]. For all experiments, the number of local training epochs was set to $r=5$, number of communication rounds was fixed at $T=100$, and the batch size was 32. The Adam optimizer has been used while the learning rate, for both global and local updates, was $lr=0.001$ across all communication rounds. The average personalized test accuracy across individual client’s data in the final communication round is reported in Table 2 and 1. Figure 3 presents the training accuracy versus communication rounds for CIFAR10 and CIFAR100 datasets.
Model Architectures: Following the FL literature, we utilized several model architectures. For MNIST, we used a simple CNN consisting of 2 convolutional layers (each with 1 block) and 2 fully connected layers. for CIFAR10 and CIFAR100, We employed a CNN with 4 convolutional layers (1 block each) and 1 fully connected layer for all datasets. Additionally, we used MobileNet, which comprises 14 convolutional layers (2 blocks each) and 1 fully connected layer, for CIFAR-10 and CIFAR-100. For partial model approaches such as FedAlt and FedSim, the split layer is fixed in the middle of the network: for CNNs, layers [1,2] are shared, while for MobileNet, layers [1–7] are considered as shared part. More details about the training process is discussed in Appendix 10.1. Our implementation is available as a supplement for reprodcing the results.
Figure 3: test accuracy accuracy curve along with global communication rounds for pMixFed and PFL baselines experimented on CIFAR10 and CIFAR100 where N=100, C=10%.
5.2 Comparative Results
We evaluate two versions of our proposed method: pMixFed, which utilizes both adaptive and dynamic layer-wise mixup degree updates, and pMixFed-Dynamic, where the mixup coefficient $\mu_{i}$ is fixed across all training rounds and gradually decreases from $1\rightarrow 0$ from the head to the base layer ($i=M\rightarrow 1$) following a Sigmoid function designed for each client. Our results show that the final accuracy of pMixFed outperforms baseline methods in most cases. As shown in Figure 3, the accuracy curve of pMixFed exhibits smoother and faster convergence, which may suggest the potential for early stopping in FL settings. Previous studies on mixup also suggest that linear interpolation between features provides the most benefit in the early training phase [65]. Moreover, Figure 3 highlights that partial models with a hard split are highly sensitive to hyperparameter selection and different distribution settings, which can lead to training instability. This issue appears to be addressed more effectively in pMixFed, as discussed further in Section 4. According to the results in Table 2 and Table 1, Ditto demonstrates relatively robust performance across different heterogeneity settings. However, its effectiveness diminishes as the model depth increases, such as with MobileNet, and under larger client populations ($N=100$). On the other hand, while FedAlt and FedSim report above-baseline results, they consistently fail during training. Adjusting hyperparameters did not resolve this issue.
5.3 Analytic Experiments
Figure 4: (a) The accuracy drop in FedSim occurred due to the vanishing gradient at round 42. (b) accuracy declines at round 10 in FedAlt due to the introduction of 5 new participants. Applying adaptive mixup solely between corresponding global and local shared layers mitigates the accuracy drop.
Adaptive Robustness to Performance Degradation: During our experiments, we observed the algorithm’s ability to adapt and recover from performance degradation, especially in challenging scenarios such as gradient vanishing, adding new users, or introducing unseen incoming data. For instance, in complex settings with larger models, such as MobileNet on the CIFAR-100 dataset, partial PFL models like FedSim and FedAlt experience sudden accuracy drops due to zero gradients or the incorporation of new participants into the cohort. We believe the reason behind this phenomenon is due to the: 1-local and global model update discrepancy in partial models with strict cut in the middle depicted in Figure 1. This degradation is mitigated by the adaptive mixup coefficient, which dynamically adjusts the degree of personalization based on the local model’s performance during both the broadcasting and aggregation stages. Specifically, if the global model $G^{(t)}$ lacks sufficient strength, the mixup coefficient $\mu^{(t)}$ is reduced, decreasing the influence of global model. 2- Catastrophic forgetting which is addressed in pMixFed by keeping the historical models $H\big{|}_{1}^{T}G^{t}$ in the aggregation process as discussed in section 4.2.2. Figure 4 illustrates that even applying mixup only to the shared layers of the same partial PFL models (FedAlt and Fedsim) enhances resilience against sudden accuracy drops, maintaining model performance over time. In both experiments, the mixup degree for personalized layers is set to $\lambda_{i}=0$ for all clients similar to FedSim and FedAlt algorithms.
5.4 Ablation Study
Random vs gradual mix factor from $\beta$ distribution: In this paper, we have explored different designs for calculating g mix factor $\mu_{k}$. The value of $\lambda$ in Eq. 4, naturally sampled from a $\beta\neq(\alpha,\alpha)$ distribution [61] which is on the interval [0,1]. We have also experimented the random $\lambda_{i}$ using $\beta$ distribution with different $\alpha$. If $\alpha=1$, the $\beta$ distribution is uniform meaning that the $\lambda$ would be sampled uniformly from [0,1]. Moreover $\alpha>1$, The $\lambda$ would be more in between, creating a more mixed output between $L_{k}$ and $G$. On the contrary, if $\alpha<1$ the mixed model tend to choose just one of the global and local models where $\lambda=1or\lambda=0$. The effects of different $\alpha$ on mixup degree $\lambda$ is discussed in Appendix.
Figure 5: (a) Effect on learning rate on average test accuracy(out-of-sample) gap and on the cold-start users. (b) The comparison between test accuracy on the cold-start-users with different Mix factor functions. (Dynamic-only). In this scenario we used a fixed $Mu$ for all communication rounds. (Sigmoid). The original updating strategy based on a sigmoid function. (Gradual). A simple linear function has been adapted for updating $Mu$. (Random) Mixup degree $\lambda_{i}$ is selected randomly from $\beta$ distribution.
Mix Factor($\mu$): sigmoid vs Dynamic-only In this study, we have exploited two different functions to update adaptive mixup factor($\mu$) in each communication round. This idea is based on the performance of the model which we want to update. In 1st scenario $\sigma$ function has been adapted as shown in equation 5 for adaptively updating mixup degree $\lambda_{i}$ using sigmoid function. On the other hand, the 2nd scenario, Dynamic-only, $Mu$ is fixed over all communication rounds. The comparison of these two scenarios as well as the effect of different $t$ values on the test accuracy, is depicted in Figure 5 (b).
Effect of Mixup Degree as Learning Rate (lr): We observed that the effect of the mixup coefficient is highly influenced by the learning rate and its decay. To empirically demonstrate this relationship, we measured the impact of the learning rate on new participants (cold-start users) as well as on the out-of-sample gap (average test accuracy on unseen data). The results of this comparison are presented in Figure 5 (a). Additional details are provided in the Appendix 10.3.
6 Conclusions
We introduced pMixFed, a dynamic, layer-wise personalized federated learning approach that uses mixup to integrate the shared global and personalized local models. Our approach features adaptive partitioning between shared and personalized layers, along with a gradual transition for personalization, enabling seamless adaptation for local clients, improved generalization across clients, and reduced risk of catastrophic forgetting. We provided a theoretical analysis of pMixFed to study the properties of its convergence. Our experiments on three datasets demonstrated its superior performance over existing PFL methods. Empirically, pMixFed exhibited faster training times, increased robustness, and better handling of data heterogeneity compared to state-of-the-art PFL models. Future research directions include exploring multi-Modal personalization and adapting pMixFed for working on resource-constrained devices
7 Acknowledgment
This work is based upon the work partly supported by the National Center for Transportation Cybersecurity and Resiliency (TraCR) (a U.S. Department of Transportation National University Transportation Center) headquartered at Clemson University, Clemson, South Carolina, USA. Any opinions, findings, conclusions, and recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of TraCR, and the U.S. Government assumes no liability for the contents or use thereof.
Appendix
8 More Detailed discussion on the Mix Factor
This section elaborates on two key properties of pMixFed: 1) dynamic behavior, achieved through a gradual transition in the mix degree between layers ($\lambda_{i}$), and 2) Adaptive behavior, introduced via the mix factor ($\mu$). Below, we delve into the details and formulation of each step.
8.1 Dynamic Mixup Degree
Among the various partitioning strategies for partial PFL discussed in the main paper and in [34], one of the most widely adopted techniques is assigning higher layers of local model $L_{l,k}^{t}$ to personalization while allowing the base layers $L_{g,k}^{t}$ to be shared across clients as the global model [33, 45, 4]. This design aligns with insights from the Model-Agnostic Meta-Learning (MAML) algorithm [10], which demonstrates that lower layers generally retain task-agnostic, generalized features, while the higher layers capture task-specific, personalized characteristics. Accordingly, in this work, we designate the head of the model as containing personalized information, while the base layers represent generalized information shared across clients  [^4] It should be noted that our method is fully adaptable to different partial PFL designs as well discussed in [34]..
Broadcasting: To achieve a nuanced and gradual transition in the mixing process between the global model and the local model, we define the mix degree $\lambda_{i}$ as follows. The local model’s head $L_{n}$ remains frozen, and the head of the global model is excluded from being shared with the local model. The mix degree for each layer $\lambda_{i}$ increases incrementally based on the mix factor $\mu$, such that as we move toward the base layers, the personalization impact decreases. This dynamic behavior is represented as:
$\lambda_{i}=\lambda_{i+1}+\mu,$
where $\lambda_{i}$ controls the degree of mixing at layer $i$. This process is visually illustrated in Figure 8.
Figure 6: layer-wise Dynamic transition of mixup degree $\lambda_{i}$ in Broadcasting and aggregation phase. The darker color shows the higher mixup degree ($\lambda_{i}$) for the corresponding layer $i$.
Aggregation: The aggregation stage focuses on preserving the generalized information from the history of previous global models to mitigate the catastrophic forgetting problem. In contrast to the broadcasting process, the primary goal here is to retain the generalized information of the previous global model, which encapsulates the history of all prior models $H\big{|}_{1}^{t-1}{G}$. Therefore, the base layers should predominantly be shared from the global model, particularly during the first round of training, where the updated local models are still underdeveloped. Nonetheless, as we transition to the head, it’s less prominent to transfer knowledge from the previous global models. Figure 8 illustrates how the mixup degree transitions from the head to the base layers.
8.2 Adaptive Mix-Factor
To enable online, adaptive updates of the mix degree $\lambda_{i}$, we implemented several algorithms, as detailed in Section 5.4 Ablation Study. We observed a clear relationship between the mixup degree and the similar behavior to lr, which is further discussed in Section 9.4.1. Inspired by learning rate schedulers, we introduced the mix-factor $\mu$ to adaptively update the layer-wise mixup degree based on the current communication round $\delta=t/T$ and the relative performance of the current local model $L_{k}^{t}$ compared to the global model $G^{t}$. The Sigmoid function in Figure 8 illustrates how $\mu$ evolves with respect to $\delta$ and accuracy ($Acc$). The best results were achieved when setting $b=1$ and using the square of $Acc$ as an exponent. The rationale behind this approach is that a more experienced, better-performing model should share more information. Specifically, if the local model accuracy $Acc_{l}$ significantly exceeds that of the global model ($Acc_{G}$) in the current round ($t$) such that $Acc\gg 1$, less information is shared from the global model. Conversely, when $10<Acc<$ 1, the global model dominates the parameter updates in both the global and local models. Figure 7 shows the distribution of the calculated $\mu$ across different communication rounds for client 0 in the broadcasting stage. In the broadcasting stage, higher $Acc_{l}$ values result in freezing more layers for personalization, leading to a decrease in $\mu$. Conversely, during the aggregation phase, if the global model accuracy $Acc_{G}$ outperforms the updated local model accuracy $Acc_{l^{\prime}}^{(t+1)}$, more base layers are shared. This is especially important during the first round of training, where local models are less stable. As demonstrated in Section 5.3, our proposed method adapts dynamically to performance drops and high-distribution complexity, adjusting the mixup degree as needed. This adaptability is effective even when applied partially to partial PFL methods such as FedAlt, and FedSim.
8.3 Model Heterogeneity
pMixFed is capable of handling variable model sizes across different clients. The global model, $M_{g}$, retains the maximum number of layers from all clients, i.e., $M_{G}=\max(M_{1},M_{2},\dots,M_{N})$. During the matching process between the global model $G_{i}$ and the local model $L_{i}$, if a layer block from the local model does not match a corresponding global layer, we set $\lambda_{i}=0$, meaning that the layer block will neither participate in the broadcasting nor aggregation processes. So the layers are participating according to their existing participation rate. For instance, if only 40% of clients have more than 4 layers, the generalization degree (in both broadcasting and aggregation stages), will be less than $0.4$ in each training round due to different participation rates.
Figure 7: mix-factors for client 0 in different communication rounds
9 Theoretical Analysis
In this section, we provide the convergence analysis of pMixFed. Moreover, We compare the aggregation process and the server global model update of the FedSGD algorithm with our proposed mixed aggregation stage in pMixFed. We begin by introducing the key notations and assumptions used throughout the convergence analysis.
Notations:
  • $t\in\{0,\ldots,T-1\}$: communication round index.
  • $\eta_{l}$: learning rate for local update, $\eta_{g}$: learning rate for global updates.
  • $\lambda_{k,i}$: mixup coefficient for client $k$ at layer $i$ in round $t$, $\lambda_{k}$: mixup coefficient for client $k$ in round $t$ (assuming uniform across layers).
  • $G^{(t)}$: global model parameters at round $t$, $L_{k}^{(t)}$: local model parameters of client $k$ at round $t$.
  • $|\mathcal{D}_{k}|$: size of the dataset at client $k$, $|\mathcal{D}|$: total size of datasets across all clients.
  • $\nabla L_{k}^{(t+1)}$: gradient of the local model at client $k$ in round $t$.
We make the following assumptions to establish the convergence properties of pMixFed:
Assumption 1.
Unbiased Gradient Estimation with Bounded Variance. The gradient estimate for local model updates is unbiased and has a bounded variance, i.e.,
$\mathbb{E}[\nabla\mathcal{L}_{k}^{(t)}]=\nabla\mathcal{L}_{k}^{(t)},\quad\text{Var}(\nabla\mathcal{L}_{k}^{(t)})\leq\sigma^{2}.$ (9)
Assumption 2 (Smoothness of Local Objectives).
The local objective functions $L_{k}(\cdot)$ are $L$-smooth, i.e.,
$\small\|\nabla L_{k}^{(t+1)}-\nabla L_{k}^{(t)}\|\leq L\|L_{k}^{(t+1)}-L_{k}^{(t)}\|.$ (10)
Assumption 3 (Bounded Gradients).
The gradients at each client are bounded, i.e.,
$\small\|\nabla L_{k}^{(t+1)}\|\leq G,\quad\forall k,t.$ (11)
These assumptions are standard in convergence analysis and ensure that the optimization process is well-behaved.
9.1 Local Training Convergence
We first analyze the local training updates that occur between communication rounds.
Lemma 1.
Local Model Training Progress. Under Assumptions 2 and 1, after $r$ local updates, the expected loss for client $k$ satisfies:
$\mathbb{E}[\mathcal{L}_{k}^{(t+r)}]\leq\mathcal{L}_{k}^{(t)}+\left(\frac{L_{1}\eta_{l}^{2}}{2}-\eta_{l}\right)\sum_{j=0}^{r-1}\|\nabla\mathcal{L}_{k}^{(t+j)}\|^{2}+\frac{L_{1}r\eta_{l}^{2}\sigma^{2}}{2}.$ (12)
This lemma provides a bound on how the local training process improves the loss function. The bound depends on the learning rate $\eta_{l}$, the smoothness constant $L_{1}$, and the variance of the gradient estimates $\sigma^{2}$.
Figure 8: Adaptive mix-factor $\mu_{k}^{t}$ according to the accuracy ration $Acc^{t}=Acc^{t}_{L}/Acc^{t}_{G}$ in different communication rounds.
9.2 Global Model Aggregation
Next, we consider the effect of aggregating local models at the server after each communication round.
Lemma 2.
Global Model Aggregation Dynamics. After aggregating the local models, the global model loss changes as follows:
$\mathbb{E}[\mathcal{L}^{(t+1)}]\leq\mathbb{E}[\mathcal{L}^{(t)}]+\eta_{g}\delta^{2},$ (13)
where $\delta^{2}$ bounds the difference between the global and local model parameters before and after aggregation.
This lemma shows that, during aggregation, the global loss increases by a term proportional to the global learning rate $\eta_{g}$ and the parameter variation $\delta^{2}$.
9.3 Convergence of pMixFed
With the above lemmas, we can now derive the convergence properties of pMixFed.
Theorem 1.
Convergence Rate of pMixFed. Under the previously stated assumptions, the expected global loss after a complete round of communication and local training satisfies:
$\small\mathbb{E}[\mathcal{L}^{(t+1)}]\leq\mathcal{L}^{(t)}+\left(\frac{L_{1}\eta_{l}^{2}}{2}-\eta_{l}\right)\sum_{k=1}^{K}\sum_{j=0}^{r-1}\|\nabla\mathcal{L}_{k}^{(t+j)}\|^{2}+\frac{L_{1}r\eta_{l}^{2}\sigma^{2}}{2}+\eta_{g}\delta^{2}.$ (14)
This theorem demonstrates that the global loss decreases with each communication round, with the convergence depending on the local and global learning rates, the gradient norms, and the parameter variation.
Theorem 2.
Non-Convex Convergence Rate. For non-convex loss functions, pMixFed achieves convergence at the following rate:
$\frac{1}{T}\sum_{t=0}^{T-1}\sum_{k=1}^{K}\|\nabla\mathcal{L}_{k}^{(t)}\|^{2}\leq\mathcal{O}(1/T),$ (15)
where $T$ is the total number of communication rounds.
This final theorem indicates that pMixFed achieves non-convex convergence with a rate of $\mathcal{O}(1/T)$, demonstrating that the algorithm improves over time as the number of communication rounds increases.
9.4 The effect of Aggregating model parameters and gradients on catastrophic forgetting
In FedSGD, the gradients are aggregated and the server will be update the global model according to the aggregated gradients. the FedSGD is sometimes preferred over FedAvg due to its potentially faster convergence. However, it lacks robustness in heterogeneous environments. pMixFed leverages the faster convergence characteristics of FedSGD by incorporating early stopping mechanisms, facilitated by the use of mixup. As demonstrated in Section 5, the mixup factor $\lambda$ functions analogously to an SGD update at the server, even though pMixFed aggregates model weights rather than gradients, similar to FedAvg. Section 9.4 provides a more detailed explanation of this mechanism.
the pMixFed algorithm combines the advantages of FedSGD and FedAvg by aggregating model weights rather than gradients, while still ensuring convergence even in heterogeneous data settings. The incorporation of the mixup mechanism enhances stability, providing faster convergence rates compared to FedSGD, particularly in non-convex settings. Considering local SGD steps $r$ and $\eta_{l}$ as the local learning rate we can further expand the above formulation:
$\begin{split}&G^{t+1}=G^{t}-\eta_{g}\sum_{i=1}^{k}\Omega_{k}\left[L_{k}^{(t+r)}-L_{k}^{(t)}\right]\\ &G^{t+1}=G^{t}-\eta_{g}.\eta_{l}\sum_{i=1}^{k}\Omega_{k}\sum_{j=0}^{r-1}\nabla{L_{k}^{(t+j)}}\end{split}$ (16)
Since in FedSGD, the global model $G^{(t)}$ is fully shared with each local model $L_{k}^{(t)}$ in each communication round t, and $\sum_{i=1}^{k}\Omega_{k}=1$, we can update the problem as a constrained optimization problem.
$G^{t+1}=G^{t}-\eta_{g}.\eta_{l}\ \sum_{j=1}^{r-1}\nabla{G^{(t+j)}}$
$\sum_{j=2}^{r-1}\nabla{G^{(t+j)}=(1-\eta_{g}.\eta_{l})G^{t}-(1+\eta_{g}.\eta_{l})G^{t+1}\ }$
The right term of the equation is the history of global model gradients which could be written as $H\big{|}_{2}^{r-1}\nabla{G^{t}}=a.G^{t}-b.G^{t+1}$. The problem of the FedSGD algorithm is that the gradients are too small to keep the state of the previous global model, leading to catastrophic forgetting in partial models that use gradients in their aggregation stage. We hypothesize that the mixup coefficient $\lambda$ acts similarly to the learning rate $\eta$, suggesting that $\lambda$ plays a role analogous to $\eta$ in FedSGD.
FedSGD Update Rule : In FedSGD, the global model $G^{(t)}$ is updated by aggregating the gradients from all clients:
$\small G^{(t+1)}=G^{(t)}-\eta_{g}\sum_{k=1}^{K}\Omega_{k}\nabla L_{k}^{(t+1)},$ (17)
where $\Omega_{k}=\frac{|\mathcal{D}_{k}|}{|\mathcal{D}|}$ is the weight associated with client $k$. Since $L_{k}^{(t+1)}=L_{k}^{(t)}-\eta_{l}\nabla L_{k}^{(t+1)}$, we can rewrite the update as:
$\small G^{(t+1)}=G^{(t)}-\eta_{g}\sum_{k=1}^{K}\Omega_{k}\left(\frac{L_{k}^{(t)}-L_{k}^{(t+1)}}{\eta_{l}}\right).$ (18)
pMixFed Update Rule: In pMixFed, the global model update incorporates mixup coefficients:
$\small G^{(t+1)}=\sum_{k=1}^{K}\Omega_{k}\left[(1-\lambda_{k})L_{k}^{(t+1)}+\lambda_{k}G^{(t)}\right].$ (19)
Assuming $\lambda_{k,i}=\lambda_{k}$ for all layers $i$, and considering that $G^{(t)}$ is sent to all clients at round $t$, we simplify the update to:
$\small G^{(t+1)}=(1-\lambda_{k})\sum_{k=1}^{K}\Omega_{k}L_{k}^{(t+1)}+\lambda_{k}G^{(t)}.$ (20)
9.4.1 Analytical analysis of the Effect of learning rate and mixup degree
To establish the relationship between $\lambda$ and $\eta_{g}$, we align the FedSGD and pMixFed update equations. From Equation (20), rearranged:
$\small G^{(t+1)}-\lambda_{k}G^{(t)}=(1-\lambda_{k})\sum_{k=1}^{K}\Omega_{k}L_{k}^{(t+1)}.$ (21)
From Equation (18), rearranged:
$\small G^{(t+1)}=G^{(t)}-\frac{\eta_{g}}{\eta_{l}}\sum_{k=1}^{K}\Omega_{k}\left(L_{k}^{(t)}-L_{k}^{(t+1)}\right).$ (22)
Assuming $L_{k}^{(t)}$ is replaced with $G^{(t)}$ at round $t$ for FedSGD, we have $L_{k}^{(t)}=G^{(t)}$. Substituting this into Equation (22):
$G^{(t+1)}=G^{(t)}-\frac{\eta_{g}}{\eta_{l}}\sum_{k=1}^{K}\Omega_{k}\left(G^{(t)}-L_{k}^{(t+1)}\right).$ (23)
Simplifying and Comparing with Equation (20), we see that if: $\lambda_{k}=\frac{\eta_{g}}{\eta_{l}}$. then the updates are analogous.
$G^{(t+1)}=G^{(t)}\left(1-\frac{\eta_{g}}{\eta_{l}}\sum_{k=1}^{K}\Omega_{k}\right)+\frac{\eta_{g}}{\eta_{l}}\sum_{k=1}^{K}\Omega_{k}L_{k}^{(t+1)}.$ (24)
Theorem 3.
Under Assumptions 2 and 3, the mixup coefficient $\lambda_{k}$ in pMixFed acts similarly to the learning rate ratio $\frac{\eta_{g}}{\eta_{l}}$ in FedSGD, such that:$\lambda_{k}=\frac{\eta_{g}}{\eta_{l}}$. This implies that the mixup mechanism in pMixFed can be interpreted as a form of learning rate control analogous to FedSGD.
Proof.
Starting from the pMixFed update in Equation (21):
$G^{(t+1)}-\lambda_{k}G^{(t)}=(1-\lambda_{k})\sum_{k=1}^{K}\Omega_{k}L_{k}^{(t+1)}.$ (25)
From the rearranged FedSGD update in Equation (24):
$G^{(t+1)}=G^{(t)}\left(1-\frac{\eta_{g}}{\eta_{l}}\right)+\frac{\eta_{g}}{\eta_{l}}\sum_{k=1}^{K}\Omega_{k}L_{k}^{(t+1)}.$ (26)
Setting the two expressions equal:
$G^{(t)}\left(1-\lambda_{k}\right)+(1-\lambda_{k})\sum_{k=1}^{K}\Omega_{k}L_{k}^{(t+1)}=$
$G^{(t)}\left(1-\frac{\eta_{g}}{\eta_{l}}\right)+\frac{\eta_{g}}{\eta_{l}}\sum_{k=1}^{K}\Omega_{k}L_{k}^{(t+1)}.$ (27)
This simplifies to:
$\lambda_{k}=\frac{\eta_{g}}{\eta_{l}}.$ (28)
Given that $\lambda_{k}$ must lie in the range $[0,1]$, this relationship holds when $\eta_{g}\leq\eta_{l}$. Since both $\eta_{g}$ and $\eta_{l}$ are also constrained within the interval $[0,1]$, our findings are consistent with previous studies [33, 10], which recommend keeping the head unfrozen (i.e., $\eta_{l}\neq 0$).
Remark1. Theorem 3 establishes a direct relationship between the mixup coefficient $\lambda_{k}$ and the learning rates used in local and global updates. This insight allows us to interpret the mixup mechanism in pMixFed as adjusting the effective learning rate at the server, providing a theoretical foundation for selecting $\lambda_{k}$ based on desired convergence properties.
Remark2. practice, this relationship suggests that by tuning $\lambda_{k}$, we can control the influence of the global model versus the local models in the aggregation process, similar to adjusting the learning rate in FedSGD. This is particularly beneficial in heterogeneous environments where clients may have varying data distributions.
Our theoretical analysis indicates that the mixup coefficient $\lambda_{k}$ in pMixFed plays a role analogous to the learning rate in FedSGD. This equivalence provides a deeper understanding of how pMixFed leverages the strengths of FedSGD while mitigating its weaknesses in heterogeneous settings. By appropriately choosing $\lambda_{k}$, pMixFed can achieve faster convergence and improved robustness.
10 Additional Details about the Experiments
10.1 Experimental Setup
Dataset: We used three widely used federated learning datasets: MNIST [23], CIFAR-10, CIFAR-100 [3], and EMNIST [8]. CIFAR-10 consists of 50,000 images of size $32\times 32$ for training and 10,000 images for testing. CIFAR-100, on the other hand, consists of 100 classes, with 500 $32\times 32$ images per class for training and 100 images per class for testing. MNIST contains 10 labels and includes 60,000 samples of $28\times 28$ grayscale images for training and 10,000 for testing. For creating heterogenity, we followed the
Training Details: For evaluation, We have reported the average test accuracy of the global model [59] for different approaches. The final global model at the last communication round is saved and used during the evaluation. The global model is then personalized according to each baseline’s personalization or fine-tuning algorithm for $r=4$ local epochs and $T=50$. For FedAlt, the local model is reconstructed from the global model and fine-tuned on the test data. For FedSim, both the global and local models are fine-tuned partially but simultaneously. In the case of FedBABU, the head (fully connected layers) remains frozen during local training, while the body is updated. Since we could not directly apply pFedHN in our platform setting, we adapted their method using the same hyper parameters discussed above and employed hidden layers with 100 units for the hypernetwork and 16 kernels. The local update process for LG-FedAvg, FedAvg, and Per-FedAvg simply involves updating all layers jointly during the fine-tuning process. The global learning rate for the methods that need sgd update in the global server e.g., FedAvg, has been set from $lr_{global}=[1e-3,1e-4,and1e-5]$.It should be noted that due to the performance drop for some methods (FedAlt , FedSim) in round 10 or 40 in some settings, we’ve reported the highest accuracy achieved. Also this is the reason the accuracy curves are illustrated for 39 rounds instead of 50.  [^5] As discussed in the main paper, the change of hyper-parameters such as lr, batch size, momentum and even changing the optimizer to adam didn’t help with the performance drop in most cases.
10.2 Out-of-Sample and Participation Gap
In the evaluation of the effect of learning rate and mixup, the average test accuracy [^6] Classification accuracy using softmax is measured on cold-start clients $|D^{ts}_{k\cap\text{unseen}}|,k\subset\{1,\dots,M\}$, where $D^{ts}\neq D^{tr}$. These clients have not participated in the federation at any point during training. FedAlt and FedSim perform poorly on cold-start users or unseen clients, highlighting their limited generalization capability. The test accuracy of pFedMix, while affected under a 10% participation rate, benefits significantly from seeing more clients. Increased client participation directly improves accuracy, as observed in previous studies [34].
10.3 Ablation Study : The Effects of Different alpha on Mixup Degree
The Beta distribution $\beta(\alpha,\alpha)$ is defined on the interval [0,1], where the parameter $\alpha$ controls the shape of the distribution. The value of $\lambda$, used in Eq. 4, is naturally sampled from this distribution. By varying the parameter $\alpha$, we can adjust how much mixing occurs between the global model $G$ and the local model $L_{k}$.
  • Uniform Distribution: When $\alpha=1$, the Beta distribution becomes uniform over [0,1]. In this case, $\lambda$ is sampled uniformly across the entire interval, meaning that each model, $G$ and $L_{k}$, has an equal probability of being weighted more or less in the mixup process. This leads to a broad exploration of different combinations of global and local models, allowing for a wide range of mixed models.
  • Concentrated Mixup ( $\alpha>1$): When $\alpha>1$, the Beta distribution is concentrated around the center of the interval [0,1]. As a result, the mixup factor $\lambda$ is more likely to be closer to 0.5, leading to more balanced combinations of the global and local models. This results in outputs that are more ”mixed,” with neither model dominating the mixup process. Such a setup can enhance the robustness of the combined model, as it prevents extreme weighting of either model, creating smoother interpolations between them.
  • Extremal Mixup ( $\alpha<1$): In contrast, when $\alpha<1$, the Beta distribution becomes U-shaped, with more probability mass near 0 and 1. This means that $\lambda$ tends to be either very close to 0 or very close to 1, favoring one model over the other in the mixup process. When $\lambda\approx 0$, the local model $L_{k}$ is chosen almost exclusively, and when $\lambda\approx 1$, the global model $G$ is predominantly selected. This form of mixup creates a more deterministic selection between global and local models, with less mixing occurring.
Figure 9: PDF $\lambda$ (mixup degree) for different values of $\alpha$ in $\beta$ distribution
The behavior of different $\alpha$ values is depicted in Figure 9, where the distribution of the mixup factor $\lambda$ is visualized. These distributions highlight the varying degrees of mixup, ranging from uniform blending to nearly deterministic model selection. To thoroughly investigate the impact of $\alpha$ on model performance, we designed two distinct experimental setups:
  • Random Sampling: In this scenario, we set different $\alpha$, meaning that the $\lambda$ values are sampled uniformly from the interval [0,1] according. This ensures a wide range of mixup combinations between the global and local models. The random sampling approach helps us assess the general robustness of the model when the mixup degree $\lambda$ is not biased towards any specific value. Table 3 shows the effect of different $\alpha$ on the overall test accuracy of pMixFed.
  • Adaptive Sampling: For this case, we divided the communication rounds into three distinct stages, each consisting of $\frac{epoch_{global}}{3}$ epochs. During these stages, the parameter $\alpha$ is adaptively changed as follows:
    $\alpha=\begin{cases}0.1,&\text{initial stage (early training)}\\ 100,&\text{middle stage (convergence phase)}\\ 10,&\text{final stage (fine-tuning)}\end{cases}$ (29)
    This adaptive strategy mimics the behavior of the original pFedMix algorithm while also allowing for more controlled exploration of different mixup combinations. During the early and late stages, a small $\alpha$ (0.1) encourages more deterministic model selections (i.e., either local or global), while the middle stage with $\alpha=100$ promotes more balanced mixing. This dynamic adjustment of $\alpha$ enables us to control the degree of mixup at different phases of training. Table 4 also shows the effect of different sampling approaches (random and adaptive) on the overall test accuracy of all three dataset pMixFed. The results shows that the adaptive sampling which creats a form scheduling for mixup degree shows promising result even compared to the original algorithm using adaptive $\mu$.
Table 3: Accuracy of $pMixFed$ with Different $\alpha$ Values in the Beta Distribution
Dataset $\alpha=0.1$ $\alpha=0.5$ $\alpha=1$ $\alpha=2$ $\alpha=5$
CIFAR-10 78.5 81.2 82.6 84.3 83.1
CIFAR-100 42.3 45.6 47.2 49.8 48.1
Table 4: Accuracy of $pMixFed$ with Different $\alpha$ Values and Sampling Strategies
Dataset Sampling Strategy Accuracy(%)
CIFAR-10 Random Sampling ($\alpha=1$) 82.6
CIFAR-10 Adaptive Sampling 85.3
CIFAR-100 Random Sampling ($\alpha=1$) 47.2
CIFAR-100 Adaptive Sampling 50.1
MNIST Random Sampling ($\alpha=1$) 97.9
MNIST Adaptive Sampling 98.6

### v2

@@
 Extensive experiments demonstrate that pMixFed outperforms state-of-the-art PFL methods, showing faster model training, increased robustness, and improved handling of data heterogeneity under different heterogeneous settings.
-Our code is available for reproducing our results: https://github.com/YasMinSdt/pMixFed.
+Keywords: Personalization, Federated Learning, Meta Learning
 1 Introduction
@@
 (Bottom) During the joint training of generalized and personalized models, the gradient updates from the generalized layers are impacted by the gradients from personalized layers, resulting in catastrophic forgetting, performance drop and slower convergence rates.
-The goal in federated learning (FL) [22] is to facilitate collaborative learning of several machine learning (ML) models in a decentralized scheme.
+AI has achieved remarkable success across computer vision, natural language processing, and healthcare tasks by leveraging large-scale datasets.
+However, in many real-world scenarios data is distributed across multiple devices or organizations, and privacy ma2020safeguarding; mothukuri2021federated; stansecure; zhu2024enhancing, regulatory cai2024esvfl; pati2024privacy; nananukul2024multi; yazdinejad2024robust, or communication constraints yao2018two; luping2019cmfl; paragliola2022definition prevent centralizing the data.
+This challenge increases heterogeneity in data distributions across clients which can degrade the performance of centralized AI methods.
+Moreover, user-specific preferences or contexts often require personalized models rather than a single global solution.
+These practical necessities motivate the need for collaborative learning methods that allow training models jointly across clients while respecting privacy, reducing communication costs, and adapting to heterogeneous clients.
+One goal of federated learning (FL) konevcny2016federated is to facilitate collaborative learning of several machine learning (ML) models in a decentralized scheme.
 FL requires addressing data privacy, catastrophic forgetting, and client drift problem  [^1] A phenomenon where the global model fails to serve as an accurate representation because local models gradually drift apart due to high data heterogeneity.
-[38, 15, 43, 31, 36].
-Existing FL methods cannot address all these challenges with non-IID (non-Independent and Identically Distributed) data.
-For instance, although “FedAvg”  [32] demonstrates strong generalization performance, it fails to provide personalized solutions for a cohort of clients with non-IID datasets.
-Hence, the global model, or one “average client” in “FedAvg”, may not adequately represent all individual local models in non-IID settings due to client-drift [53].
+rostami2018multi; huang2022learn; singhal2021federated; luo2023gradma; qu2022rethinking.
+Existing FL methods cannot address all these challenges with non-Independent and Identically Distributed (non-IID) data.
+For instance, although Federated Averaging (“FedAvg”) mcmahan2017communication demonstrates strong generalization performance, it fails to provide personalized solutions for a cohort of clients with non-IID datasets.
+Hence, the global model, or one “average client” in “FedAvg”, may not adequately represent all individual local models in non-IID settings due to client-drift xiao2020averaging.
 Personalized FL (PFL) methods handle data heterogeneity by considering both generalization and personalization during the training stage.
-Since there is a trade-off between generalization and personalization in heterogeneous environments, PFL methods leverage heterogeneity and diversity as advantages rather than adversities [35, 47].
+Since there is a trade-off between generalization and personalization in heterogeneous environments, PFL methods leverage heterogeneity and diversity as advantages rather than adversities pye2021personalised; tan2022towards.
 A group of PFL approaches train personalized local models on each device while collaborating toward a shared global model.
 Partial PFL, also known as parameter decoupling, involves using a partial model sharing, where only a subset of the model is shared while other parameters remain “frozen” to balance generalization and personalization until the subsequent round of local training.
-While partial PFL methods are effective in mitigating catastrophic forgetting, strengthening privacy, and reducing computation and communication overhead [34, 45], there are still some unaddressed challenges.
+While partial PFL methods are effective in mitigating catastrophic forgetting, strengthening privacy, and reducing computation and communication overhead pillutla2022federated; sun2023fedperfix, there are still some unaddressed challenges.
 First, the question of when, where, and how to optimally partition the full model?
 remains unresolved.
-Recent studies [34, 45] have shown that there is no “one-size-fits-all” solution; the best or optimal partitioning strategy depends on factors such as task type (e.g., next word prediction or speech recognition) and local model architecture.
+Recent studies pillutla2022federated; sun2023fedperfix have shown that there is no “one-size-fits-all” solution; the best or optimal partitioning strategy depends on factors such as task type (e.g., next word prediction or speech recognition) and local model architecture.
 An improper partitioning choice can lead to issues such as underfitting, overfitting, increased bias, and catastrophic forgetting.
-Some studies [29] suggest that personalized layers should reside in the base layers, while others [9, 4] argue that the base layers contain more generalized information and should be shared.
+Some studies liang2020think suggest that personalized layers should reside in the base layers, while others collins2021exploiting; arivazhagan2019federated argue that the base layers contain more generalized information and should be shared.
 Further, the use of a fixed partitioning strategy across all communication rounds for heterogeneous clients can limit the efficacy of collaborative learning.
 For instance, if the performance of the client suddenly drops due to new incoming data, the partitioning strategy should be changed because the client requires more frozen layers.
-Another issue is catastrophic forgetting of the previously shared global knowledge after only a few rounds of local training because the shared global model can be completely overwritten by local updates leading to generalization degradation [31, 41, 15, 54].
+Another issue is catastrophic forgetting of the previously shared global knowledge after only a few rounds of local training because the shared global model can be completely overwritten by local updates leading to generalization degradation luo2023gradma; shirvani2023comprehensive; huang2022learn; xu2022acceleration.
 Most importantly, partial models may experience slower convergence compared to full model personalization, as frozen local model updates can diverge in an opposite direction from the globally-shared model.
 Since the generalized and personalized models are trained on non-IID datasets, there might also be a domain shift, leading to model discrepancy as depicted in Figure 1.
-These discrepancies arise from variations in local and global objective functions, differences in initialization, and asynchronous updates [56, 24].
-As a result, merging the shared and the personalized layers can disrupt information flow within the network, impede the learning process, and lead to a slower convergence rate or accuracy drop in partial PFL models such as FedAlt and FedSim [34] [^2] More details on this is discussed in section 4..
-Further, while partial PFL techniques contribute to an overall improved training accuracy, they can reduce the test accuracy on some devices, particularly in devices with limited samples, leading to variations in results in terms of the performance level [34].
+These discrepancies arise from variations in local and global objective functions, differences in initialization, and asynchronous updates yang2024fedas; lee2023partial.
+As a result, merging the shared and the personalized layers can disrupt information flow within the network, impede the learning process, and lead to a slower convergence rate or accuracy drop in partial PFL models such as FedAlt and FedSim pillutla2022federated [^2] More details on this is discussed in section 4..
+Further, while partial PFL techniques contribute to an overall improved training accuracy, they can reduce the test accuracy on some devices, particularly in devices with limited samples, leading to variations in results in terms of the performance level pillutla2022federated.
 Hence, there is a need for novel solutions to achieve the following goals in PFL:
   • Dynamic and Adaptive Partitioning: The balance between shared and personalized layers should be dynamically and adaptively adjusted for each client during every communication round, rather than relying on a static, fixed partitioning strategy for all participants.
-  • Gradual Personalization Transition: The degree of personalization should transition gradually across layers, as opposed to an ”all-or-nothing” approach that employs strict partitioning or hard splits within the model discussed in Figure 1.
+  • Gradual Personalization Transition: The degree of personalization should transition gradually across layers, as opposed to an “all-or-nothing” approach that employs strict partitioning or hard splits within the model discussed in Figure 1.
 This ability allows nuanced adaptation for individual client needs due to heterogeneity.
@@
   • Scalability and Adaptability: The approach should be fast, scalable, and easily adaptable to new cold-start clients while accounting for model/device heterogeneity.
-To achieve the above, we propose “pMixFed”, a layer-wise, dynamic PFL approach that integrates Mixup [61] between the shared global and personalized local models’ layers during both the broadcasting (global model sharing with local clients) and aggregation (aggregating distributed local models to update the global model) stages within a partial PFL framework.
+To achieve the above, we propose “pMixFed”, a layer-wise, dynamic PFL approach that integrates Mixup zhang2017mixup between the shared global and personalized local models’ layers during both the broadcasting (global model sharing with local clients) and aggregation (aggregating distributed local models to update the global model) stages within a partial PFL framework.
 Our main contributions include:
-  • We develop an online, dynamic interpolation method between local and global models using Mixup [57], effectively addressing data heterogeneity and scalable across varying cohort sizes, degrees of data heterogeneity, and diverse model sizes and architectures.
+  • We develop an online, dynamic interpolation method between local and global models using Mixup yoon2021fedmix, effectively addressing data heterogeneity and scalable across varying cohort sizes, degrees of data heterogeneity, and diverse model sizes and architectures.
   • Our solutions facilitate a gradual increase in the degree of personalization across layers, rather than relying on a strict cut-off layer, helping to mitigate client drift.
@@
   • “pMixFed” reduces the participant gap (test accuracy for cold-start users) and the out-of-sample gap (test accuracy on unseen data) caused by data heterogeneity through linear interpolation between client updates, thereby mitigating the impact of client drift.
+The remainder of this paper is organized as follows.
+In Section 2, we review the relevant literature on personalized federated learning, highlighting existing methodologies and their limitations.
+Section 3 introduces the problem formulation, including both fully personalized and partially personalized federated learning settings, accompanied by theoretical analysis.
+In Section 5, we present the proposed pMixFed framework, detailing the integration of adaptive and dynamic mixup strategies across different model layers.
+Section 6 provides extensive empirical evaluations of pMixFed, including comparisons with state-of-the-art baselines and analyses of computational and communication efficiency.
+Finally, Section 7 concludes the paper and outlines potential directions for future research.
 2 Related Work
-PFL aims to adapt local models to the individual needs, preferences, and contexts of each participant.
-Since the inception of FL, various PFL approaches have been explored to address the challenges stemming from heterogeneity, categorized into three areas: 1) data-centric strategies 2) global model adaptation, and 3) local model personalization.
-The following provide a more detailed discussion of these three approaches:
-Data-centric: These approaches in PFL aim to address data heterogeneity and class imbalances by manipulating statistical data properties, such as data size, distribution, and local data selection [47].
-Examples of these techniques include data normalization, feature engineering, data augmentation, synthetic data generation, and client selection.
-Representative methods, such as Astraea [11], P2P k-SMOTE [50], FedMCCS [1], FedAug [19], [63], and FedHome [52], exemplify data-centric approaches.
-The shortcoming of these methods is that they often modify the natural distribution and statistical characteristics of federated data, potentially injecting bias or eliminating valuable, rare information.
-Additionally, many of these strategies depend on the availability of proxy data on the global server, which may conflict with the privacy regulations of some entities.
-Global Model Adaptation: In these approaches, a single global model is maintained on the server and subsequently adapted to individual local models in the following phase.
-The primary goals of these methods are: 1) learning a robust, generalized model and 2) enabling fast, efficient local model adaptation.
-Several techniques employ regularization terms to mitigate client drift and prevent model divergence during local updates e.g., FedProx [27], SCAFFOLD [21], FL-MOON [26], and FedCurv [42].
-Some of these approaches such as PerFedAvg [12], pFedMe [46], and pFedHN [39] leverage Meta-learning, and transfer learning e.g., FedStag [55], and DPFed [58], yet they introduce specific challenges.
-Meta-learning methods can be computationally intensive, while regularization techniques may add computational overhead by incorporating additional terms in the objective function.
-Similarly, transfer learning algorithms can be inefficient in terms of communication overhead and often require a public dataset to enhance the global model on the server.
-A common limitation across most adaptation techniques is the need for uniform model architecture across all clients, requiring devices with varying computational capabilities to use the same model size.
-Local Model Personalization: The limitations of existing PFL methods have led to the development of local model personalization approaches that focus on training customized models for each client.
-Some approaches utilize multi-task learning (MTL), a collaborative framework that facilitates information exchange across distinct tasks (e.g., FedRes [2], MOCHA [44], FedAMP [16], and Ditto [28]).
-Another category leverages knowledge distillation (KD) to support personalization when client-specific training objectives differ, as seen in FedDF [30], FedMD [25], FedGen [64], and FedGKT [14].
-However, both MTL and KD can incur significant computational and communication overhead, limiting their scalability for large-scale FL deployments across resource-constrained devices.
-Partial PFL: These methods, also known as parameter decoupling, are the primary focus of this paper.
-These methods mitigate catastrophic forgetting by retaining shared components while freezing other layers for personalization.
-FedPer [4] introduced partial models in FL, sharing only initial layers with generalized information and reserving final layers for personalization.
-FedBABU [33] divides the network into a shared body and a frozen head with fully connected layers.
-Other frameworks, such as FURL [6] and LG-FedAvg [29], apply partial PFL by retaining private feature embeddings or using compact representation learning for high-level features, respectively.
-Two baseline methods in this paper are FedAlt [43] and FedSim [34].
+PFL seeks to adapt each client’s model to its individual data, preferences, and context.
+Similar to classic domain adaptation, data privacy can be an important concern in PFL stan2024unsupervised; verma2025leveraging; bingtao2025prodg; tian2024novel, the additional challenge, however, is that data is distributed and we need to adapt several models due to data heterogeneity between the clients.
+Since the advent of FL, a wide range of PFL methods have been proposed to address client heterogeneity (both statistical and system), which we group into four categories: (i) data-centric strategies; (ii) clustering-based approaches; (iii) global-model adaptation; and (iv) local-model personalization—the focus of this work, which also encompasses partial-model (parameter-decoupling) techniques.
+We discuss each category in turn below.
+2.1 Data-Centric
+Data-centric methods in PFL seek to mitigate distribution shift and class imbalance by shaping the data seen during training.
+for example, by adjusting sample sizes, label distributions, and client selection strategies tan2022towards.
+Common techniques include data normalization, feature engineering, augmentation, synthetic data generation, and adaptive client sampling.
+Representative examples include Astraea duan2019astraea, which performs reputation-based client selection by evaluating neighbors’ updates and rewarding trustworthy contributors; P2P k-SMOTE wang2021non, where clients synthesize minority-class examples via $k$-nearest-neighbor interpolation to reduce local imbalance; and FedMCCS 9212434, which improves fairness by selecting participants based on multiple factors (e.g., dataset size, class diversity, and training loss) rather than at random.
+FedAug jeong2018communication; zhao2018federated employs server-side data augmentation to alleviate heterogeneity, while FedHome wu2020fedhome augments local data using autoencoders trained with server-provided synthetic samples to address scarcity and imbalance.
+More recent works directly target class imbalance in PFL: Lee et al.
+lee2024regularizing construct a separate global model per class and let clients personalize by composing these class-wise globals according to local distributions; Liang et al.
+liang2024fedrema (FedReMa) selectively aggregates updates from the most relevant clients to enhance personalization under skewed data; and FedGA xiao2024fedga aligns client gradients before aggregation to reduce imbalance-induced bias and improve minority-class performance.
+Despite their promise, data-centric approaches alter the underlying statistics of federated data and may inadvertently introduce bias or erase rare but important patterns.
+Several methods also rely on proxy or auxiliary server-side data, raising privacy and distribution-mismatch concerns.
+Finally, the added computation (and sometimes communication) can be burdensome for resource-constrained devices, limiting the practicality of these strategies in real-world PFL deployments.
+2.2 Clustering
+Clustering-based methods in PFL address the core challenge of heterogeneity: each client may hold data from markedly different distributions (e.g., distinct users, sensors, or hospitals).
+Naïvely averaging updates (e.g., FedAvg) can induce negative transfer when distributions diverge.
+Instead of a single global model, these approaches exploit structure across clients so that those with similar distributions reinforce one another while avoiding distortion from dissimilar updates.
+Clients thus retain the benefits of collaboration while mitigating negative transfer; in large populations, clustering/graph constructions also reduce complexity relative to fully personalized models.
+Clustering methods can be grouped into three categories: (1) Dynamic Clustering and Graph-based Aggregation; (2) Per-Client Weighting; and (3) Personalized Clustering.
+In the first category, methods dynamically group clients or build graphs to personalize aggregation based on inter-client relationships.
+For example, FEDCEDAR wang2024rethinking builds dynamic weighted graphs among clients, enabling precise, personalized model distribution via dynamic graph propagation.
+Multi-Center FL long2023multi learns multiple global model “centers,” assigning each client to its best-fitting center through an EM-style procedure, thereby improving personalization under heterogeneous distributions.
+More recently, song2025chpfl proposes a three-layer clustered hierarchical PFL framework that adapts clustering at multiple levels to handle severe non-IID data.
+In the second category (Per-Client Weighting), instead of uniform averaging, methods assign personalized aggregation weights according to client similarity or utility.
+FedSPD lin2024fedspd uses soft clustering so that clients converge toward cluster-relevant models while preserving personalization through flexible model consensus.
+FedDWA liu2023feddwa employs a parameter server to compute dynamic, per-client aggregation weights from model-update similarity, reducing communication and privacy overhead.
+In the third category (Personalized Clustering), personalization is organized around classifier outputs or by merging multiple personalized models.
+For instance, pFedCk zhang2024personalized combines clustering with knowledge distillation: each client keeps a personalized model locally while an interaction model is shared on the server; features and logits are distilled across similar clients (clustered by the server) to enhance personalization.
+Likewise, COMET cho2023communication clusters clients based on distribution similarity and enables knowledge transfer through logits-based co-distillation, yielding cluster-specific personalized models while maintaining model heterogeneity and communication efficiency.
+Clustering can be computationally heavy, and assignments may be unstable—clients can be misassigned when data are noisy or highly imbalanced.
+Maintaining multiple cluster models becomes costly at scale.
+Finally, similarity measures or EM responsibilities may leak distribution information, especially in centroid-based or hierarchical clustering.
+2.3 Global Model Adaptation:
+In these approaches, a single global model is maintained on the server and subsequently adapted to individual local models in a later phase.
+The primary goals are: (1) learning a robust, generalized model and (2) enabling fast, efficient local adaptation.
+Several techniques employ regularization terms to mitigate client drift and prevent model divergence during local updates.
+For example, FedProx li2020federated adds a proximal term so local updates stay close to the current global; SCAFFOLD karimireddy2020scaffold uses control variates to cancel client drift while training a single server model; MOON li2021model introduces a contrastive term to keep local representations aligned with the global and/or the previous local model; and FedCurv shoham2019overcoming infuses an EWC-style curvature penalty into the objective to keep local steps near a shared optimum.
+Another line of single-global PFL approaches leverages meta-learning and transfer learning.
+Methods such as PerFedAvg fallah2020personalized learn a global initialization that each client can adapt with a few steps (MAML in FL).
+In FedL2P lee2023fedl2p meta-nets learn client-specific personalization hyperparameters (e.g., layerwise LRs / BN) to fine-tune the same global efficiently.
+MetaVers lim2024metavers meta-learns a versatile representation so the shared model adapts rapidly at clients.
+FedSteg yang2020fedsteg uses federated transfer/knowledge-transfer ideas for image steganalysis, and DPFed yu2020salvaging explicitly studies fine-tuning/local adaptation of a federated global model.
+Global model adaptation introduces specific challenges.
+Meta-learning methods can be computationally intensive, while regularization-based techniques add overhead by incorporating extra terms into the objective.
+Similarly, transfer-learning approaches can be communication-inefficient and often require a public dataset to enhance the server-side global model.
+A common limitation across most adaptation techniques is the need for a uniform model architecture across all clients, forcing devices with varying computational capabilities to use the same model size.
+Another relevant direction is zero-shot learning (ZSL), which aims to train a model that generalizes on unseen classes, tasks, or domains without having direct training data for them hao2021towards; zhang2022fedzkt; rostami2022zero; asif2024advanced; guo2025feze.
+In the context of federated learning, ZSL can be interpreted as enabling the global or personalized models to perform well for new clients or unseen data distributions without requiring additional local training rounds.
+Recent works have leveraged generative models asif2024advanced to bridge the gap between seen and unseen tasks.
+Integrating ZSL capability into PFL could improve support for cold-start clients by allowing them to benefit from global knowledge without significant local updates.
+2.4 Local Model Personalization:
+The limitations of existing PFL methods have led to approaches that train customized models for each client.
+One line of work utilizes multi-task learning (MTL), a collaborative framework that facilitates information exchange across distinct tasks.
+for example, MOCHA smith2017federated, a canonical federated MTL method that trains related but distinct client models with a shared regularizer; and FedAMP huang2021personalized, where attentive message passing builds client-specific models with pairwise collaboration.
+In Ditto li2021ditto, each client optimizes its own personalized model with a regularization term.
+FEDHCA2 lu2024fedhca2 learns relationships between heterogeneous clients to produce personalized but collaborative models.
+In FedRes agarwal2020federated, clients learn local residuals that add to a shared global model; the personalized local predictor (global + local) can be viewed through the lens of MTL or parameter decoupling which is fully discussed below.
+Another category leverages knowledge distillation (KD) to support personalization when client-specific training objectives differ.
+In FedMD li2019fedmd, clients keep heterogeneous local models and distill via public logits to learn personalized local models.
+In FedGen zhu2021data, a server-side generator helps clients personalize via KD and feature transfer.
+FedGKT he2020group provides group teacher–student training and supports client-side students to form personalized models.
+These KD-based methods often include an adaptive fusion step that balances generalization and personalization without requiring full model sharing.
+Some approaches also combine KD with other categories; for example, pFedCK zhang2024personalized combines client clustering and KD to produce personalized client models, and FedGMKD zhang2024fedgmkd uses prototype-based KD with client-specific distillation to target local personalization.
+A third, fast-growing line is hypernetwork or weight-generation–based personalization.
+Although the number of works is still relatively small, the space is expanding with adapters, prompts, and personalized parameter generation.
+The first paper, pFedHN shamsian2021personalized, trains a server-side hypernetwork that generates client-specific parameters from client embeddings; each client then receives its own weights, and personalization requires minimal client training.
+HyperFLoRA lu2024hyperflora trains a hypernetwork that generates LoRA adapter modules personalized to each client for LLM personalization: clients send lightweight statistics, and the server generates per-client adapters.
+Both MTL- and KD-based approaches can incur significant computational and communication overhead, limiting scalability for large-scale FL deployments on resource-constrained devices.
+Many KD-based PFL methods assume access to proxy/public datasets for logit matching or generator training, which is unrealistic in privacy-sensitive settings.
+Hypernetworks, on the other hand, face a capacity bottleneck: a single hypernetwork must generate useful weights for many diverse clients; under high heterogeneity, it may underfit or collapse toward “average” solutions.
+A prominent subcategory of local-model personalization, which decouples a shared parameter set (e.g., backbone) from client-specific modules (e.g., heads/adapters), is called partial models.
+This design is widely adopted because it is parameter-efficient, easy to quantize or deploy, and operationally transparent.
+By keeping the shared component stable and personalizing only a small subset, these methods reduce negative transfer across clients and mitigate catastrophic forgetting of global knowledge while enabling fast on-device adaptation.
+FedPer arivazhagan2019federated introduced partial models in FL, sharing only initial layers with generalized information and reserving final layers for personalization.
+FedBABU oh2021fedbabu divides the network into a shared body and a frozen head with fully connected layers.
+Other frameworks, such as FURL bui2019federated and LG-FedAvg liang2020think, apply partial PFL by retaining private feature embeddings or using compact representation learning for high-level features, respectively.
+Two baseline methods in this paper are FedAlt singhal2021federated and FedSim pillutla2022federated.
 FedAlt uses a stateless FL paradigm, reconstructing local models from the global model, while FedSim synchronously updates shared and local models with each iteration.
-Limitations: These methods face challenges such as model update discrepancies (as shown in Figure 1) and catastrophic forgetting, where shared layers may undergo significant changes after only a few rounds of local training, resulting in sudden accuracy drops.
+FedSelect tamirisa2024fedselect is a Lottery-ticket-inspired subnetwork selection method that has been published recently.
+This method gradually personalizes a subnetwork per client while aggregating the rest.
+FedTSDP zhu2024federated is a PFL method that combines partial-model with clustering to better handle non-IID data.
+FDLoRA qi2024fdlora, which is an LLM-based FL systems, use LoRA (Low-Rank Adaptation) technique to improve personalization.
+FDLoRA introduces dual LoRA modules per client, one capturing global knowledge aggregated by the server, and the other capturing client-specific personalization.
+An adaptive fusion process reconciles the two, effectively balancing generalization with personalization without requiring full model sharing.
+These methods face challenges such as model update discrepancies (as shown in Figure 1) and catastrophic forgetting, where shared layers may undergo significant changes after only a few rounds of local training, resulting in sudden accuracy drops especially in high-scale Fl setting.
 Additionally, users with high personalization accuracy may freeze more layers than cold-start users or unreliable participants, who should rely more heavily on the global model.
@@
 $F_{k}(\theta)=\frac{1}{|\mathcal{D}_{k}|}\sum_{(x,y_{i})\in\mathcal{D}_{k}}\ell(f_{k}(\theta,x_{i}),y_{i})$
-where $|\mathcal{D}|=\sum_{k=1}^{K}|\mathcal{D}_{k}|$ and $F(\theta)$ is the global loss function of FedAvg [32].
+where $|\mathcal{D}|=\sum_{k=1}^{K}|\mathcal{D}_{k}|$ and $F(\theta)$ is the global loss function of FedAvg mcmahan2017communication.
 FL is performed in at iterative fashion.
@@
 In Equation 1, the assumption is that the data is collected from an IID distribution, and all clients should train their data according to the exact same model.
-However, this assumption cannot be applied within many practical FL settings due to the non-IID nature of data and resource limitations [18, 17].
+However, this assumption cannot be applied within many practical FL settings due to the non-IID nature of data and resource limitations imteaj2021survey; imteaj2021fedparl.
 In the PFL settings with high heterogeneity and non-IID data distribution, the same issue persists and the local parameters need to be customized toward each agent.
-PFL extends FL by solving the following [26, 34]:
+PFL extends FL by solving the following li2021model; pillutla2022federated:
 $\min_{\theta,\theta_{k}}\sum_{k=1}^{K}\frac{1}{|\mathcal{D}_{k}|}(\mathcal{F}_{k}(\theta_{k})+\alpha_{k}\|\mathbf{\theta_{k}-\theta}\|^{2}).$ (2)
@@
 When $\alpha_{k}$ is small, the personalization power of the local models will be increased.
-If $\alpha_{k}$ is large, the local models’ parameters tend to be closer to the global model parameters $\theta$.
+If $\alpha_{k}$ is large, the local models’ parameters tend to be closer to the global parameters $\theta$.
 3.1 Partial Personalized Model
@@
 $\mathcal{F}_{k}(\theta_{k})=\ell(f_{k}(\langle L_{g,k};L_{l,k}\rangle,x_{k}),y_{k})$ (3)
-Among different partitioning strategies for partial PFL [34], the most popular technique is to assign local personalized layers $L_{l,k}^{t}$ to final layers and allow the base layers $L_{g,k}^{t}$ to share the knowledge similar to FedPer [4].
+Among different partitioning strategies for partial PFL pillutla2022federated, the most popular technique is to assign local personalized layers $L_{l,k}^{t}$ to final layers and allow the base layers $L_{g,k}^{t}$ to share the knowledge similar to FedPer arivazhagan2019federated.
 This choice aligns with insights from MAML  [^3] Model-Agnostic Meta-Learning algorithm, suggesting that initial layers keeps general and broad information while personalized characteristics manifest prominently in the higher layers.
@@
 In the next broadcasting phase, the shared layers of the local model will be updated as $L_{g}^{(t+1)}\leftarrow G^{(t+1)}$.
-Our goal is to benefit from Mixup to improve personalization.
-4 The Proposed Method
+Table 1: Notation used throughout the paper. Vectors are column vectors; $\|\cdot\|$ is the Euclidean norm.
+Symbol Type Meaning
+$K$ scalar Number of clients
+$\mathcal{U}_{t}$ set Clients selected at round $t$; $|\mathcal{U}_{t}|{=}K$
+$D_{k}$ dataset Local dataset of client $k$; size $|D_{k}|$
+$|D|$ scalar Total data size, $|D|{=}\sum_{k}|D_{k}|$
+$\Omega_{k}$ scalar Aggregation weight $|D_{k}|/|D|$
+$\ell(\cdot,\cdot)$ function Pointwise loss used to build local/global objectives
+$f_{k}(\theta,x)$ function Model prediction on client $k$ with parameters $\theta$
+$F_{k}(\theta)$ function Local objective on client $k$
+$F(\theta)$ function Global objective $\sum_{k}\Omega_{k}F_{k}(\theta)$
+$\theta$ vector Global (server) model parameters
+$\theta_{k}$ vector Client $k$’s local/personal parameters (if stored)
+$G^{(t)}$ vector Server/global parameters at round $t$ (alias of $\theta^{t}$)
+$L_{k}^{(t)}$ vector Local model of client $k$ at round $t$
+$L^{(t)}_{g,k}$ vector Global/shared layers of $L_{k}^{(t)}$
+$L^{(t)}_{l,k}$ vector Local/personalized layers of $L_{k}^{(t)}$
+$M$ scalar Number of layers in the network
+$s$ index Split (cut) layer index, $1\leq s\leq M$
+$t$ index Communication round index ($t{=}0,\dots,T{-}1$)
+$T$ scalar Total number of communication rounds
+$r,\ \tau$ scalar Number of local steps per round (local iterations/epochs)
+$b$ scalar Local mini-batch size
+$\eta_{\ell}$ scalar Local learning rate (client step size)
+$\eta_{g}$ scalar Server/aggregation step size in the FedSGD view
+$\lambda^{t}_{k,i}$ scalar Mixup coefficient for client $k$, layer $i$, round $t$ ($\in[0,1]$)
+$\lambda^{t}_{k}$ scalar Client-level mixup (uniform across layers if assumed)
+$\bar{\lambda}^{\,t}$ scalar Population-averaged mixup, $\sum_{k\in\mathcal{U}_{t}}\Omega_{k}\lambda_{k}^{t}$
+$\eta_{\mathrm{eff}}^{\,t}$ scalar Effective step size $(1{-}\bar{\lambda}^{\,t})\,\eta_{\ell}$ in SGD view
+$\mu$ scalar Mix factor controlling the schedule of $\lambda$ across layers/rounds
+$\nabla F_{k}(\theta)$ vector Full local gradient on client $k$
+$g_{k}^{t}$ vector Stochastic gradient on client $k$ at round $t$
+$g_{k,s}^{t}$ vector Stochastic gradient at local step $s\in\{0,\dots,\tau{-}1\}$ in round $t$
+$L$ scalar Smoothness constant (Lipschitz gradient)
+$\mu_{\text{sc}}$ scalar Strong convexity constant (when assumed in Sec. 4)
+$\kappa$ scalar Condition number $L/\mu_{\text{sc}}$ (strongly convex case)
+$\sigma^{2}$ scalar Upper bound on gradient variance
+$\zeta_{k}$ scalar Gradient dissimilarity: $\sup_{\theta}\|\nabla F(\theta){-}\nabla F_{k}(\theta)\|^{2}$
+$\zeta$ scalar Aggregate dissimilarity (e.g., $\sum_{k}\zeta_{k}$ or mean variant)
+$\Delta_{k}$ scalar Local–global optimum gap $\|\theta_{k}^{\star}{-}\theta^{\star}\|^{2}$
+$\mathcal{G}$ scalar Gradient moment bound used in multi-step analysis (Sec. 4)
+$\mathrm{Beta}(\alpha_{\mathrm{B}},\alpha_{\mathrm{B}})$ dist. Beta distribution used to sample $\lambda$ in MixUp
+Notes.
+(i) Aggregation-as-FedSGD yields $\boxed{\bar{\lambda}^{\,t}=1-\eta_{g}/\eta_{\ell}}$ and $\eta_{\mathrm{eff}}^{\,t}=(1{-}\bar{\lambda}^{\,t})\eta_{\ell}$.
+(ii) We reserve $\mu_{\text{sc}}$ for strong convexity; $\mu$ denotes the mixup scheduling factor if present in Methodology.
+4 Theoretical Analysis
+In this section we establish a clean theoretical view of pMixFed.
+We show that the update rule is equivalent to stochastic gradient descent (SGD) on the global objective with an effective step size governed by the mixup coefficients.
+This both clarifies stability conditions and explains how pMixFed mitigates catastrophic forgetting.
+In FedSGD, the gradients are aggregated and the server will be update the global model according to the aggregated gradients.
+FedSGD is sometimes preferred over FedAvg due to its potentially faster convergence.
+However, it lacks robustness in heterogeneous environments.
+pMixFed leverages the faster convergence characteristics of FedSGD by incorporating early stopping mechanisms, facilitated by the use of mixup.
+As It’ll be demonstrated in Section 5, the mixup factor $\lambda$ functions analogously to an SGD update at the server, even though pMixFed aggregates model weights rather than gradients, similar to FedAvg.
+Section 4.2 provides a more detailed explanation of this mechanism.
+4.1 Assumptions
+Let $F(\theta):=\sum_{k=1}^{K}\Omega_{k}F_{k}(\theta)$ denote the global objective, where $\Omega_{k}=|D_{k}|/|D|$ and $\theta^{t}$ is the global model at round $t$.
+Each client $k$ runs $r$ local SGD steps with step size $\eta_{\ell}$, producing $\theta_{k}^{t+1}$.
+Mixup coefficients $\lambda_{k,i}^{t}\in[0,1]$ apply at layer $i$ of client $k$.
+For simplicity we define client-averaged $\lambda_{k}^{t}:=\sum_{i}w_{i}\lambda_{k,i}^{t}$ with $\sum_{i}w_{i}=1$, and population average $\bar{\lambda}^{t}:=\sum_{k}\Omega_{k}\lambda_{k}^{t}$.
+We make the following standard assumptions.
+These assumptions are standard in convergence analysis and ensure that the optimization process is well-behaved.
+Assumption 1
+(Unbiased Gradient Estimation with Bounded Variance.) The gradient estimate for local model updates is unbiased and has a bounded variance, i.e.,)
+$\mathbb{E}[g_{k}^{t}]=\nabla F_{k}(\theta^{t})\quad\text{and}\quad\mathbb{E}\|g_{k}^{t}-\nabla F_{k}(\theta^{t})\|^{2}\leq\sigma^{2}.$ (4)
+Assumption 2 (L-Smoothness)
+Each $F_{k}$ has $L$-Lipschitz gradients, i.e.,
+$\|\nabla F_{k}(x)-\nabla F_{k}(y)\|\leq L\|x-y\|$ (5)
+Assumption 3 (Bounded Gradients)
+The gradients at each client are bounded, i.e.,
+$\mathbb{E}\|g_{k}^{t}\|^{2}\leq G^{2}$ (6)
+4.2 Equivalence to FedSGD
+In pMixFed, the server forms a convex combination between the previous global iterate and the locally updated models:
+$\theta^{t+1}\penalty 10000\ =\penalty 10000\ \sum_{k\in\mathcal{U}_{t}}\Omega_{k}\left(\lambda_{k}^{t}\,\theta^{t}+(1-\lambda_{k}^{t})\,\theta_{k}^{t+1}\right)\penalty 10000\ =\penalty 10000\ \bar{\lambda}^{\,t}\,\theta^{t}+(1-\bar{\lambda}^{\,t})\sum_{k\in\mathcal{U}_{t}}\Omega_{k}\theta_{k}^{t+1}.$ (7)
+With one local step from $\theta^{t}$, $\theta_{k}^{t+1}=\theta^{t}-\eta_{\ell}\,g_{k}^{t}$, (7) becomes
+$\theta^{t+1}\penalty 10000\ =\penalty 10000\ \theta^{t}-\underbrace{(1-\bar{\lambda}^{\,t})\,\eta_{\ell}}_{\displaystyle\eta_{\mathrm{eff}}^{\,t}}\sum_{k\in\mathcal{U}_{t}}\Omega_{k}\,g_{k}^{t}.$ (8)
+Thus pMixFed is exactly centralized SGD on $F$ with effective step size
+$\boxed{\qquad\eta_{\mathrm{eff}}^{\,t}\penalty 10000\ =\penalty 10000\ (1-\bar{\lambda}^{\,t})\,\eta_{\ell}.\qquad}$
+Matching the FedSGD view.
+If the server instead writes a FedSGD step with global step size $\eta_{g}>0$, i.e., $\theta^{t+1}=\theta^{t}-\eta_{g}\sum_{k}\Omega_{k}g_{k}^{t}$, then (8) shows the coefficients must satisfy:
+Theorem 1 (Coefficient matching with FedSGD)
+For any round $t$, the pMixFed update (7) similar with a FedSGD step of size $\eta_{g}$ if and only if
+$\boxed{\qquad\bar{\lambda}^{\,t}\penalty 10000\ =\penalty 10000\ 1-\frac{\eta_{g}}{\eta_{\ell}}\,,\qquad 0\leq\frac{\eta_{g}}{\eta_{\ell}}\leq 1.\qquad}$
+Proof Equate (8) with $\theta^{t+1}=\theta^{t}-\eta_{g}\sum_{k}\Omega_{k}g_{k}^{t}$ and match the coefficient of the aggregated gradient: $(1-\bar{\lambda}^{\,t})\eta_{\ell}=\eta_{g}$.
+Interpretation.
+Larger $\eta_{g}/\eta_{\ell}$ implies smaller $\bar{\lambda}^{\,t}$, i.e., less carry-over of $\theta^{t}$ and more trust in the new local updates.
+Standard smoothness suggests choosing $\eta_{\mathrm{eff}}^{\,t}\leq 1/L$, equivalently $\bar{\lambda}^{\,t}\geq 1-\frac{1}{L\eta_{\ell}}$ whenever $L\eta_{\ell}>1$.
+4.3 Convergence guarantees
+The SGD view (8) directly yields standard optimization guarantees. We first state a one-step descent inequality, then derive the nonconvex and strongly-convex global results.
+Lemma 2 (One-step descent)
+Under (Asuumption1)–(Assumption2) and $\eta_{\mathrm{eff}}^{\,t}\leq 1/L$,
+$\mathbb{E}\big[F(\theta^{t+1})\big]\penalty 10000\ \leq\penalty 10000\ \mathbb{E}\big[F(\theta^{t})\big]\penalty 10000\ -\penalty 10000\ \frac{\eta_{\mathrm{eff}}^{\,t}}{2}\,\mathbb{E}\big\|\nabla F(\theta^{t})\big\|^{2}\penalty 10000\ +\penalty 10000\ \frac{L\,(\eta_{\mathrm{eff}}^{\,t})^{2}}{2}\,\sigma^{2}.$
+Proof By $L$-smoothness, $F(\theta^{t+1})\leq F(\theta^{t})+\langle\nabla F(\theta^{t}),\theta^{t+1}-\theta^{t}\rangle+\frac{L}{2}\|\theta^{t+1}-\theta^{t}\|^{2}$.
+Substitute (8), take conditional expectation using $\mathbb{E}[\sum_{k}\Omega_{k}g_{k}^{t}]=\nabla F(\theta^{t})$ and $\mathbb{E}\|\sum_{k}\Omega_{k}g_{k}^{t}\|^{2}\leq\|\nabla F(\theta^{t})\|^{2}+\sigma^{2}$, and use $\eta_{\mathrm{eff}}^{\,t}\leq 1/L$ to absorb the quadratic term in $\|\nabla F(\theta^{t})\|^{2}$.
+Theorem 3 (Nonconvex rate)
+Let $\eta_{\mathrm{eff}}^{\,t}\equiv\eta_{\mathrm{eff}}\leq 1/L$ be constant. Then for any $T\geq 1$,
+$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\big\|\nabla F(\theta^{t})\big\|^{2}\penalty 10000\ \leq\penalty 10000\ \frac{2\,(F(\theta^{0})-F^{\star})}{T\,\eta_{\mathrm{eff}}}\penalty 10000\ +\penalty 10000\ L\,\eta_{\mathrm{eff}}\,\sigma^{2}.$
+Proof Sum Lemma 2 over $t=0,\dots,T-1$ and telescope; divide by $T$. Please see appendix . 9
+Theorem 4 (Strongly-convex case)
+Under (A1)–(A3), with $\eta_{\mathrm{eff}}^{\,t}\equiv\eta_{\mathrm{eff}}\leq\min\{1/L,\,1/(2\mu)\}$,
+$\mathbb{E}\big\|\theta^{t}-\theta^{\star}\big\|^{2}\penalty 10000\ \leq\penalty 10000\ (1-\mu\,\eta_{\mathrm{eff}})^{t}\,\big\|\theta^{0}-\theta^{\star}\big\|^{2}\penalty 10000\ +\penalty 10000\ \mathcal{O}\!\left(\frac{\sigma^{2}}{\mu}\,\eta_{\mathrm{eff}}\right).$
+Proof Apply the standard strongly-convex SGD recursion with step $\eta_{\mathrm{eff}}$ to (8). Please see appendix . 9
+4.4 Layer-wise mixing and forgetting control
+Layer-wise coefficients $\lambda^{t}_{k,i}$ (with averaging weights $w_{i}$) produce the same effective step via $\bar{\lambda}^{\,t}=\sum_{k,i}\Omega_{k}\,w_{i}\,\lambda^{t}_{k,i}$ and $\eta_{\mathrm{eff}}^{\,t}=(1-\bar{\lambda}^{\,t})\eta_{\ell}$.
+The per-round drift satisfies
+$\big\|\theta^{t+1}-\theta^{t}\big\|\penalty 10000\ =\penalty 10000\ \eta_{\mathrm{eff}}^{\,t}\,\Big\|\sum_{k\in\mathcal{U}_{t}}\Omega_{k}\,g_{k}^{t}\Big\|.$
+Hence larger $\bar{\lambda}^{\,t}$ explicitly caps the amount a single round can overwrite the global representation, limiting catastrophic forgetting, while smaller $\bar{\lambda}^{\,t}$ accelerates adaptation.
+Maintaining $\eta_{\mathrm{eff}}^{\,t}\leq 1/L$ ensures stability irrespective of the layer schedule.
+4.5 Multiple local steps per round
+Suppose each selected client performs $\tau\geq 1$ local SGD steps from $\theta^{t}$: $\theta_{k}^{t+1}\penalty 10000\ =\penalty 10000\ \theta^{t}-\eta_{\ell}\sum_{s=0}^{\tau-1}g_{k,s}^{t}.$ Then
+$\theta^{t+1}\penalty 10000\ =\penalty 10000\ \theta^{t}\penalty 10000\ -\penalty 10000\ (1-\bar{\lambda}^{\,t})\,\eta_{\ell}\sum_{k\in\mathcal{U}_{t}}\Omega_{k}\sum_{s=0}^{\tau-1}g_{k,s}^{t}\penalty 10000\ \equiv\penalty 10000\ \theta^{t}\penalty 10000\ -\penalty 10000\ \eta_{\mathrm{eff},\tau}^{\,t}\,\widehat{g}^{\,t},\qquad\eta_{\mathrm{eff},\tau}^{\,t}:=(1-\bar{\lambda}^{\,t})\,\eta_{\ell}\,\tau,$
+where $\widehat{g}^{\,t}$ is the average of the $\tau$ local gradients. Under (Assumption1)–(Assumption2) and bounded second moments, standard local-SGD arguments give
+$\big\|\mathbb{E}[\widehat{g}^{\,t}]-\nabla F(\theta^{t})\big\|\penalty 10000\ \leq\penalty 10000\ c_{1}\,L\,\eta_{\ell}\,(\tau-1)\,G,$
+for a constant $c_{1}$ and gradient-moment bound $G$.
+Consequently, pMixFed behaves like SGD with step $\eta_{\mathrm{eff},\tau}^{\,t}$ up to a bias of order $(1-\bar{\lambda}^{\,t})\,L\,\eta_{\ell}^{2}\,\tau(\tau-1)$.
+Choosing $\bar{\lambda}^{\,t}$ so that $\eta_{\mathrm{eff},\tau}^{\,t}\leq 1/L$ recovers the guarantees of Theorems 34 with an additional (controlled) residual proportional to $L\,\eta_{\ell}^{2}\,\tau(\tau-1)$.
+Remark1.
+Theorem 1 establishes a direct relationship between the mixup coefficient $\lambda_{k}$ and the learning rates used in local and global updates.
+This insight allows us to interpret the mixup mechanism in pMixFed as adjusting the effective learning rate at the server, providing a theoretical foundation for selecting $\lambda_{k}$ based on desired convergence properties.
+Remark2.
+In practice, this relationship suggests that by tuning $\lambda_{k}$, we can control the influence of the global model versus the local models in the aggregation process, similar to adjusting the learning rate in FedSGD.
+This is particularly beneficial in heterogeneous environments where clients may have varying data distributions.
+Our theoretical analysis indicates that the mixup coefficient $\lambda_{k}$ in pMixFed plays a role analogous to the learning rate in FedSGD.
+This equivalence provides a deeper understanding of how pMixFed leverages the strengths of FedSGD while mitigating its weaknesses in heterogeneous settings.
+By appropriately choosing $\lambda_{k}$, pMixFed can achieve faster convergence and improved robustness.
+The pMixFed algorithm combines the advantages of FedSGD and FedAvg by aggregating model weights rather than gradients, while still ensuring faster convergence even in heterogeneous data settings due to it’s adaptation ability.
+The incorporation of the mixup mechanism enhances stability, providing faster convergence rates compared to FedSGD, particularly in non-convex settings
+5 Methodology
+In this section, we present our FL approach, which incorporates layer-wise Mixup in the feature space.
+We begin by briefly defining the Mixup technique, followed by a detailed description of the proposed algorithm and its integration into the federated setting.
+We also discuss the rationale behind this design and explain how it facilitates improved personalization across clients.
 5.1 Mixup
-Mixup is a data augmentation technique for enhancing model generalization [61] based on learning to generalize on linear combinations of training examples.
+Mixup is a data augmentation technique for enhancing model generalization zhang2017mixup based on learning to generalize on linear combinations of training examples.
 Variations of Mixup have consistently excelled in vision tasks, contributing to improved robustness, generalization, and adversarial privacy.
 Mixup creates augmented samples as:
-$\small\begin{split}\bar{x}=\lambda.x_{i}+(1-\lambda).x_{j}\\ \bar{y}=\lambda.y_{i}+(1-\lambda).y_{j},\end{split}$ (4)
+$\begin{split}\bar{x}=\lambda.x_{i}+(1-\lambda).x_{j}\\ \bar{y}=\lambda.y_{i}+(1-\lambda).y_{j},\end{split}$ (9)
 where $x_{i}$ and $x_{j}$ are two input samples, $y_{i}$ and $y_{j}$ are the corresponding labels, $\lambda$, $\lambda\sim\text{Beta}(\alpha,\alpha),\lambda\in[0,1]$, is the degree of interpolation between the two samples.
-Mixup relates data points belonging to different classes which has been shown to be successful in mitigating overfitting and improving model generalization [49, 13, 62].
+Mixup relates data points belonging to different classes which has been shown to be successful in mitigating overfitting and improving model generalization verma2019manifold; guo2019mixup; zhang2020does.
 There are many variants of Mixup that has been developed to address specific challenges and enhance its effectiveness.
-For instance, AlignMixup [48] improves local spatial alignment by introducing transformations that better preserve semantic consistency between input pairs.
-Manifold Mixup [49] extends the concept to hidden layer representations, acting as a powerful regularization technique by training deep neural networks (DNNs) on linear combinations of intermediate features.
-CutMix [60] replaces patches between images, blending visual information while retaining spatial structure.
-Remix [7] addresses class imbalance by assigning higher weights to minority classes during the mixing process, enhancing the robustness of the trained model on imbalanced datasets.
-AdaMix [13] dynamically optimizes the mixing distributions, reducing overlaps and improving training efficiency.
+For instance, AlignMixup venkataramanan2022alignmixup improves local spatial alignment by introducing transformations that better preserve semantic consistency between input pairs.
+Manifold Mixup verma2019manifold extends the concept to hidden layer representations, acting as a powerful regularization technique by training deep neural networks (DNNs) on linear combinations of intermediate features.
+CutMix yun2019cutmix replaces patches between images, blending visual information while retaining spatial structure.
+Remix chou2020remix addresses class imbalance by assigning higher weights to minority classes during the mixing process, enhancing the robustness of the trained model on imbalanced datasets.
+AdaMix guo2019mixup dynamically optimizes the mixing distributions, reducing overlaps and improving training efficiency.
 This linear interpolation also serves as a regularization technique that shapes smoother decision boundaries, thereby enhancing the ability of a trained model to generalize to unseen data.
-Mixup can also increase robustness against adversarial attacks [62, 5]; and improves performance against noise, corrupted labels, and uncertainty as it relaxes the dependency on specific information [13].
+Mixup also increases robustness against adversarial attacks zhang2020does; beckham2019adversarial; and improves performance against noise, corrupted labels, and uncertainty as it relaxes the dependency on specific information guo2019mixup.
 Figure 2: Workflow of pMixFed: Mixup is used in two stages.
 1-Broadcasting: when transferring knowledge to local models, the frozen personalized model $L_{k}^{(t)}$ is mixed up with global model $G^{(t)}$ according to the adaptive mix factor $\mu_{k}^{(t)}$ which determines layer-wise mixup degree $\lambda_{i}$ for layer $i$.
-2-Aggregation: The updated global model $G^{(t+1)}$ is generated through a Mixup between the updated local model $L^{(t+1)}$ and the current shared global model $G^{\prime(t)}$.
-Consequently, a state of the previous global model $G^{\prime(t)}$ is retained, which aids in mitigating catastrophic forgetting.
+2-Aggregation: The updated global model $G^{(t+1)}$ is generated through applying Mixup between the updated local model $L^{(t+1)}$ and the current global model $G^{\prime(t)}$ state.
 5.2 pMixFed : Partial Mixed up Personalized federated learning
 Our goal is to leverage the well-established benefits of Mixup in the context of personalized FL.
-While Mixup has previously been employed in FL frameworks, such as XORMixup [40], FEDMIX [57], and FedMix [51], prior studies have primarily focused on using Mixup for data augmentation or data averaging.
-We propose pMixFed by integrating Mixup on the model parameter space, rather using it on the feature space.
+While Mixup has previously been employed in FL frameworks, such as XORMixup shin2020xor, FEDMIX yoon2021fedmix, and FedMix wicaksana2022fedmix, prior studies have primarily focused on using Mixup for data augmentation or data averaging.
+We propose pMixFed by integrating Mixup on the model parameter space, rather than using it on the feature space.
 We apply Mixup between the parameters of the global and the local models in a layer-wise manner for more customized and adaptive PFL.
 Our approach eliminates the need for static and rigid partitioning strategies.
-Specifically, during both the broadcasting and aggregation stages of our partial PFL framework, we generate mixed model weights using a an interpolation strategy which is illustrated in Figure 2.
+Specifically, during both the broadcasting and aggregation stages of our partial PFL framework, we generate mixed model weights using an interpolation strategy which is illustrated in Figure 2.
 Mixup offers the flexibility in combining models by introducing a mix degree for each layer $\lambda_{i}$, which changes gradually according to $\mu$, i.e., the mix factor.
@@
 $\mu$ is computed as follows:
-$\small\mu_{k}^{t}=1-\frac{1}{1+e^{-\delta(Acc^{b}-1)}}$ (5)
+$\mu_{k}^{t}=1-\frac{1}{1+e^{-\delta(Acc^{b}-1)}}$ (10)
 $\delta=(t/T)$, where $t$, and $T$ are the current communication round, and the overall number of communication rounds respectively.
@@
 $b$ is the offset parameter for sigmoid function which is set to $2$ after several experiments.
-More detailed discussion on parameter $\mu$’s rule of update is discussed in section 5 and the reason behind this formulation is discussed in the Appendix 8.
+More detailed discussion on parameter $\mu$’s rule of update is discussed in section 5 and the reason behind this formulation is discussed in the Appendix Sec.
+2.
 As shown in Figure 2, Mixup is applied in two distinct stages of FL.
@@
 Calculation of the Mixup degree of layer $i$ $\lambda_{i}$ at both broadcasting and aggregation stages is performed as follows:
-$\small\begin{split}\text{Broadcasting Stage:}\quad\quad\lambda_{i}=\begin{cases}1&\lambda_{i}>1\\ \mu*(n-i)&\lambda_{i}\leq 1\end{cases}\\ \text{Aggregation Stage:}\quad\quad\lambda_{i}=\begin{cases}0&\lambda_{i}\leq 0\\ 1-(i*\mu)&\lambda_{i}>0,\end{cases}\end{split}$ (6)
+$\begin{split}\text{Broadcasting Stage:}\quad\quad\lambda_{i}=\begin{cases}1&\lambda_{i}>1\\ \mu*(n-i)&\lambda_{i}\leq 1\end{cases}\\ \text{Aggregation Stage:}\quad\quad\lambda_{i}=\begin{cases}0&\lambda_{i}\leq 0\\ 1-(i*\mu)&\lambda_{i}>0,\end{cases}\end{split}$ (11)
 where $n$ is total number of layers, $i$ is the current layer number starting from the base layer to the head as $i=0\rightarrow i=n$.
 $\mu$ is the mix factor which will be adaptively updated in each communication round $t$ and for each local model $k$ according to Eq.
-5.
-Algorithm 1 pMixFed: Broadcasting global to local model
-1:Input: Initial states global model: $GM^{(0)}$, local models: $\{LM^{(0)}_{i}\}_{i=1,\ldots,M}$, Number of communication rounds $T$, number of devices per round $m$, Number of layers in local models $\{L_{i}\}$
-2:for $t=0,1,\ldots,T-1$ do
-3:    Server selects $K$ devices $S(t)\subset\{1,\ldots,N\}$
-4:    Update $\mu$ for each $LM_{i}$, $i=\{1,\ldots,K\}$
-5:    Server broadcasts $GM^{(t)}$ to each device in $S(t)$
-6:    for each device $m\in S(t)$ in parallel do
-7:        $(LM^{(t+1)}_{m},GM^{(t+1)}_{m})=\text{Mixup}[(LM^{\prime(t+1)}_{m},GM^{\prime(t+1)}_{m}),\mu]$
-8:        Device sends $GM^{(t+1)}_{m}$ back to server
-9:        Update $\mu_{i=1,\ldots,K}$
-10:    end for
-11:    Server updates $GM^{(t+1)}=\frac{1}{K}\sum_{i\in S(t)}GM^{(t+1)}_{m}$
-12:end for
-Algorithm 2 Proposed Mixup for Aggregation
-1:Input: Initial states global model: $GM^{(0)}$, Number of communication rounds $T$, Number of local iterations $Itr$, number of devices per round $m$, Mixup degree $\lambda$
-2:for $t=0,1,\ldots,T-1$ do
-3:    Server broadcasts $GM^{(t)}$ to each device in $S(t)$: $LM_{i}^{(t)}$
-4:    Update $\lambda_{i=1,\ldots,M}$ for each device
-5:    for each device $i\in S(t)$ do
-6:        for $epoch=0,1,\ldots,Itr-1$ do
-7:           Train local model:
-8:                         $LM_{i}^{(t+1)}=GM^{(t)}$
-9:                     $GM^{(t+1)}=$
-10:$\hskip 65.44133pt\text{Mixup}(LM_{i}^{(t+1)},GM^{(t)})$
-11:        end for
-12:        Device sends $GM^{(t+1)}$ back to server
-13:        Adaptively update $\lambda_{i=1,\ldots,M}$
-14:    end for
-15:end for
+10.
 5.2.1 Broadcasting : Global to local model transfer
@@
 With this adaptive and flexible approach, not only can upcoming streaming unseen data be managed, but also the participation gap (test accuracy of new cold start users) would be improved.
-The Broadcasting phase is illustrated in 1 which is inspired by [20].
+The Broadcasting phase is illustrated in Figure 1 which is inspired by kairouz2021advances.
 The update rule of local model $L_{k}^{(t)}$ is as follows:
-$\small\begin{split}&{L^{\prime}}^{(t)}_{k,i}=(\lambda^{(t)}_{k,i}).G^{(t)}_{i}+(1-\lambda^{(t)}_{k,i}).L^{(t)}_{k,i}\\ &L_{k}^{(t+1)}={L^{\prime}}^{(t)}_{k}-\eta\nabla F_{k}({L^{\prime}}^{(t)}_{k}).\end{split}$ (7)
+[…]
 5.2.2 Aggregation: Local to Global Model Transfer
@@
 The updated global layers from different clients, $G_{k}^{(t+1)}$, are typically aggregated using Equation 3, which can lead to catastrophic forgetting.
-Since The base layers of the global model serve as the backbone of shared knowledge [37], This issue arises because, during each local update, the generalized layers undergo substantial modifications [34, 33].
+Since The base layers of the global model serve as the backbone of shared knowledge raghu2019rapid, This issue arises because, during each local update, the generalized layers undergo substantial modifications pillutla2022federated; oh2021fedbabu.
 When the global model is updated by simply aggregating local models, valuable information from previously shared knowledge may be lost, leading to forgetting—even if $G^{t}$ performs better than the newly aggregated model.
@@
 It should be noted that parameter $\mu$ is constant in the aggregation stage for all clients as it’s dependent to the average performance of the previous global model.
-$\small\begin{split}&{G^{\prime}}^{(t)}_{k,i}=(\lambda^{(t)}_{i,k}).G^{(t)}_{i}+(1-\lambda^{(t)}_{k,i}).L^{(t+1)}_{k,i},\\ &G^{(t+1)}=\sum_{k=1}^{K}\frac{|D_{k}|}{\sum_{k=1}^{K}|D_{k}|}G^{\prime(t+1)}_{k}.\end{split}$ (8)
+[…]
 The high-level block-diagram visualization of the proposed method is shown in Figure 2.
@@
 The parameter $\lambda$ in Equation 9 determines the Mixup degree between the shared model $GM$ and the local models $LM^{(\cdot)}_{i}$, while $\mu$ governs the slope of the change in $\lambda$ across different layers.
-The degree of Mixup gradually decreases according to the parameter $\mu$ from 00 to $1$.
+The degree of Mixup gradually decreases according to the parameter $\mu$ from 0 to $1$.
 In this scenario, $\lambda=0$ for the first base layer, indicating total sharing, while $\lambda=1$ applies to the final layer, which represents no sharing.
@@
 In each training round, only one client is “Mixed up” with the global model and the $\lambda$ parameter is adaptively learned based on the objective function using online learning.
-For a theoretical convergence analysis of pMixFed, please refer to the Appendix 9.
+Algorithm 1 Broadcasting: global to local model transfer
+1:Input: Initial states global model: $G^{(0)}$, local models: $\{L^{(0)}_{i}\}_{i=1,\ldots,N}$, Number of communication rounds $T$, Number of communication rounds $T$, Number of local training rounds ItrItr, Number of layers in local models $\{L_{i}\}$
+2:for $t=0,1,\ldots,T-1$ do
+3:  Server selects $K$ devices $S(t)\subset\{1,\ldots,N\}$
+4:  Update $\mu$ for each $L_{i}$, $i=\{1,\ldots,K\}$
+5:  Server broadcasts $G^{(t)}$ to each device in $S(t)$
+6:  for each device $k\in S(t)$ do
+7:   $L^{\prime(t)}_{k}=\text{Mix}[(L^{(t)}_{k},G^{(t)}_{k}),\mu_{Broad}]$
+8:   for $l=0,1,\ldots,Itr-1$ do
+9:     $L^{\prime(t+1)}_{k}\xleftarrow{Localtrain}(L^{\prime(t)}_{k},\eta)$
+10:   end for
+11:   Each device saves the updated model $L^{(t+1)}_{k}=L^{\prime(t+1)}_{k}$
+12:  end for
+13:end for
+Algorithm 2 Aggregation:local to global model transfer
+1:Input: Initial states global model: $G^{(0)}$, Number of communication rounds $T$, number of devices per round $s(t)$, Mix factor $\mu$
+2:for $t=0,1,\ldots,T-1$ do
+3:  for each device $k\in S(t)$ do
+4:   Update $\mu^{t}_{k}$ for $k$ in round $t$
+5:   $G^{\prime(t+1)}_{k}=\text{Mix}[(L^{(t+1)}_{k},G^{(t)}_{k}),\mu_{Agg}]$
+6:  end for
+7:  Each Device sends $G^{\prime(t+1)}_{k}$ back to server
+8:  $G^{(t+1)}\xleftarrow{Agg}{G^{\prime(t+1)}_{k}}_{k\in S(t)}$
+9:end for
 6 Experiments
-Table 1: Comparison of the State-of-The-Art Methods using Mobile-Net model
+Table 2: Average Test Accuracy of different methods using Mobile-Net model. See more details in section 6.2
 Method CIFAR-10 CIFAR-100
 $N=100$ $N=10$ $N=100$ $N=10$
-$C=10\%$ $C=100\%$ $C=10\%$ $C=100\%$ $C=10\%$ $C=100\%$ $C=10\%$ $C=100\%$
+$\% C=100C=1$$0\%$ $\% C=100C=1$$0\%$ $\% C=100C=1$$0\%$ $\% C=100C=1$$0\%$
 FedAvg 26.93 25.61 15.64 19.55 4.50 4.54 17.65 21.24
@@
 Ditto 26.58 62.77 18.60 43.98 7.00 16.07 16.89 20.74
+FedRep 29.60 31.25 32.46 51.73 14.23 28.03 33.75 45.76
 Per-FedAvg 34.30 43.59 19.52 38.29 24.04 30.65 16.27 37.20
@@
 pMixFed 69.94 72.42 54.90 74.62 45.62 56.63 54.71 58.25
-Table 2: Comparison of the State-of-The-Art Methods using CNN model
+Table 3: Average Test Accuracy of different methods using CNN model. See more details in section 6.2
 Method CIFAR-10 CIFAR-100 MNIST
 $N=100$ $N=10$ $N=100$ $N=10$ $N=100$ $N=10$
-$C=10\%$ $C=100\%$ $C=10\%$ $C=100\%$ $C=10\%$ $C=100\%$ $C=10\%$ $C=100\%$ $C=100\%$ $C=100\%$
+$\% C=100C=1$$0\%$ $\% C=100C=1$$0\%$ $\% C=100C=1$$0\%$ $\% C=100C=1$$0\%$ $\% C=100C=100\%$
 FedAvg 54.78 56.82 44.11 54.37 25.77 26.73 34.47 39.93 97.54 98.59
@@
 FedBaBU 53.12 54.60 39.77 53.21 16.77 17.33 25.60 32.47 98.19 99.07
-Ditto 46.86 79.60 31.65 60.75 27.16 42.93 25.38 55.27 98.03 95.51
+Ditto 46.86 79.60 31.65 60.75 27.16 42.93 25.38 35.27 98.03 95.51
+FedRep 53.38 57.97 47.77 67.32 34.67 31.04 24.95 39.96 97.04 98.71
 Per-FedAvg 39.37 45.03 10.00 48.13 32.67 39.01 8.71 41.21 98.32 50.34
@@
 6.1 Experimental Setup
-Datasets: We used three datasets widely used federated learning: MNIST [23], CIFAR-10, and CIFAR-100 [3], We followed the setup in [32] and [33] to simulate heterogeneous, non-IID data distributions across clients for both train and test datasets.
+6.1.1 Datasets:
+We used three datasets widely used federated learning: MNIST lecun1998gradient, CIFAR-10, CIFAR-100 alex2009learning, and Caltech-101.
+CIFAR-10 consists of 50,000 images of size $32\times 32$ for training and 10,000 images for testing.
+CIFAR-100, on the other hand, consists of 100 classes, with 500 $32\times 32$ images per class for training and 100 images per class for testing.
+MNIST contains 10 labels and includes 60,000 samples of $28\times 28$ grayscale images for training and 10,000 for testing.
+Caltech-101 is a canonical benchmark for object recognition, introduced by Li and colleagues in 2003.
+The dataset comprises 9,146 images distributed across 101 object categories—such as helicopter, elephant, and chair—plus a background/clutter class.
+Category sizes range from roughly 40 to 800 images (with most near 50), and images are typically around 300×200 pixels.
+We followed the setup in mcmahan2017communication and oh2021fedbabu to simulate heterogeneous, non-IID data distributions across clients for both train and test datasets.
 The maximum number of classes per user is set to $S=5$ for CIFAR-10 and MNIST, and $S=50$ for CIFAR-100.
 Experiments were conducted across varying heterogeneous settings, including small-scale ($N=10$) and large-scale ($N=100$) client populations, with different client participation rates $C=[100\%,10\%]$ to measure the effects of stragglers.
-Baselines and backbone: We compare two version of our method (i) pMixFed: an adaptive and dynamic mixup-based PFL approach, and (ii) pMixFed-Dynamic: a dynamic-only mixup variant where the parameter $\mu$ is fixed across communication rounds, against several baselines.
-These baselines include: FedAvg [32], FedAlt [34], FedSim [34], FedBABU [33].
-Additionally, we compare against full model personalization methods, Ditto [28], Per-FedAvg [12] , and LG-FedAvg [29].
-For all experiments, the number of local training epochs was set to $r=5$, number of communication rounds was fixed at $T=100$, and the batch size was 32.
+For Caltech-101 experiment setup, please refer to the appendix 10.3.
+6.1.2 Training Details:
+For evaluation, we have reported the average test accuracy of the global model yuan2021we for different approaches.
+The final global model at the last communication round is saved and used during the evaluation.
+The global model is then personalized according to each baseline’s personalization or fine-tuning algorithm for $r=4$ local epochs and $T=50$.
+For FedAlt, the local model is reconstructed from the global model and fine-tuned on the test data.
+For FedSim, both the global and local models are fine-tuned partially but simultaneously.
+In the case of FedBABU, the head (fully connected layers) remains frozen during local training, while the body is updated.
+Since we could not directly apply pFedHN in our platform setting, we adapted their method using the same hyper parameters discussed above and employed hidden layers with 100 units for the hypernetwork and 16 kernels.
+The local update process for LG-FedAvg, FedAvg, and Per-FedAvg simply involves updating all layers jointly during the fine-tuning process.
+The global learning rate for the methods that need sgd update in the global server e.g., FedAvg, has been set from $lr_{global}=[1e-3,1e-4,and1e-5]$.It should be noted that due to the performance drop for some methods (FedAlt , FedSim) in round 10 or 40 in some settings, we’ve reported the highest accuracy achieved.
+Also this is the reason the accuracy curves are illustrated for 39 rounds instead of 50.
+[^4] Change of hyper-parameters such as lr, batch size, momentum and even changing the optimizer to adam didn’t help with the performance drop in most cases.
+6.1.3 Baselines and backbone:
+We compare our method pMixFed: an adaptive and dynamic mixup-based PFL approach, against several baselines.
+These baselines include: FedAvg mcmahan2017communication, FedAlt pillutla2022federated, FedSim pillutla2022federated, FedBABU oh2021fedbabu.
+Additionally, we compare against full model personalization methods, Ditto li2021ditto, FedRep yang2023fedrep Per-FedAvg fallah2020personalized , and LG-FedAvg liang2020think.
+For all experiments, the number of local training epochs was set to $r=4$, number of communication rounds was fixed at $T=50$, and the batch size was 32.
 The Adam optimizer has been used while the learning rate, for both global and local updates, was $lr=0.001$ across all communication rounds.
-The average personalized test accuracy across individual client’s data in the final communication round is reported in Table 2 and 1.
+The average personalized test accuracy across individual client’s data in the final communication round is reported in Table 3 and 2.
 Figure 3 presents the training accuracy versus communication rounds for CIFAR10 and CIFAR100 datasets.
-Model Architectures: Following the FL literature, we utilized several model architectures.
+6.1.4 Model Architectures:
+Following the FL literature, we utilized several model architectures.
 For MNIST, we used a simple CNN consisting of 2 convolutional layers (each with 1 block) and 2 fully connected layers.
@@
 For partial model approaches such as FedAlt and FedSim, the split layer is fixed in the middle of the network: for CNNs, layers [1,2] are shared, while for MobileNet, layers [1–7] are considered as shared part.
-More details about the training process is discussed in Appendix 10.1.
-Our implementation is available as a supplement for reprodcing the results.
-Figure 3: test accuracy accuracy curve along with global communication rounds for pMixFed and PFL baselines experimented on CIFAR10 and CIFAR100 where N=100, C=10%.
+More details about the training process are discussed in Appendix Sec.
+4.1.
+Experimental Setup.
+Our implementation is also available as a supplement for reproducing the results.
+Figure 3: Average Test accuracy in different global communication rounds for pMixFed and other PFL baselines experimented on CIFAR10 and CIFAR100 where N=100, C=10%.
+More details are discussed in Section 6.3
 6.2 Comparative Results
-We evaluate two versions of our proposed method: pMixFed, which utilizes both adaptive and dynamic layer-wise mixup degree updates, and pMixFed-Dynamic, where the mixup coefficient $\mu_{i}$ is fixed across all training rounds and gradually decreases from $1\rightarrow 0$ from the head to the base layer ($i=M\rightarrow 1$) following a Sigmoid function designed for each client.
-Our results show that the final accuracy of pMixFed outperforms baseline methods in most cases.
+We evaluated our proposed method: pMixFed, which utilizes both adaptive and dynamic layer-wise mixup degree updates, and pMixFed-Dynamic, where the mixup coefficient $\mu_{i}$ is fixed across all training rounds and gradually decreases from $1\rightarrow 0$ from the head to the base layer ($i=M\rightarrow 1$) following a Sigmoid function designed for each client.
+Our results show that the average test accuracy of pMixFed outperforms baseline methods in most cases.
 As shown in Figure 3, the accuracy curve of pMixFed exhibits smoother and faster convergence, which may suggest the potential for early stopping in FL settings.
-Previous studies on mixup also suggest that linear interpolation between features provides the most benefit in the early training phase [65].
+Previous studies on mixup also suggest that linear interpolation between features provides the most benefit in the early training phase zou2023benefits.
 Moreover, Figure 3 highlights that partial models with a hard split are highly sensitive to hyperparameter selection and different distribution settings, which can lead to training instability.
@@
 On the other hand, while FedAlt and FedSim report above-baseline results, they consistently fail during training.
-Adjusting hyperparameters did not resolve this issue.
+[^5] Adjusting hyperparameters did not resolve this issue.
+Caltech-101.
+To further evaluate robustness under high inter-class variability and limited per-class samples, we conduct experiments on Caltech-101, a canonical benchmark for heterogeneous federated learning.
+This dataset poses unique challenges due to its long-tail distribution (most categories have $\sim$50 images) and high intra-class variability.
+We follow the same federated setup as in CIFAR/MNIST experiments, splitting the data among $N=\{10,100\}$ clients with participation rates $C=\{10\%,100\%\}$.
+As in prior work, each client is restricted to at most $S=30$ classes.
+We compare pMixFed against two recent state-of-the-art personalization baselines: FedRep yang2023fedrep, which learns shared representations with personalized heads, and Per-FedAvg fallah2020personalized, which meta-learns a global initialization for fast personalization.
+We also include FedAlt pillutla2022federated as a representative partial-sharing baseline.
+Table 4: Caltech-101 top-1 accuracy (%) with a CNN backbone across four federated settings. Values are the average across evaluation angles A–D; full A–D breakdowns are reported in Appendix 10.3.
+Method $N{=}100,\penalty 10000\ C{=}100\%$ $N{=}10,\penalty 10000\ C{=}100\%$ $N{=}100,\penalty 10000\ C{=}10\%$ $N{=}10,\penalty 10000\ C{=}10\%$
+pMixFed (Ours) 79.3 82.4 76.1 79.0
+FedRep 77.0 80.1 74.2 76.9
+Per-FedAvg 76.5 81.0 73.5 76.2
+FedAlt 73.1 75.5 69.5 72.2
+Across all settings, pMixFed consistently achieves the highest accuracy (Table 4).
+In particular, under the most heterogeneous regime (N=100N{=}100, $C{=}10\%$), pMixFed improves by $\sim$2% over FedRep and $\sim$3% over Per-FedAvg, while FedAlt lags by 6–7%.
+Per-FedAvg performs competitively in the low-client regime ($N=10$), but its advantage diminishes as heterogeneity increases.
+These results validate that pMixFed not only excels on CIFAR and MNIST benchmarks, but also generalizes robustly to harder, imbalanced datasets like Caltech-101.
 6.3 Analytic Experiments
@@
 The value of $\lambda$ in Eq.
-4, naturally sampled from a $\beta\neq(\alpha,\alpha)$ distribution [61] which is on the interval [0,1].
+9, naturally sampled from a $\beta\neq(\alpha,\alpha)$ distribution zhang2017mixup which is on the interval [0,1].
 We have also experimented the random $\lambda_{i}$ using $\beta$ distribution with different $\alpha$.
@@
 On the contrary, if $\alpha<1$ the mixed model tend to choose just one of the global and local models where $\lambda=1or\lambda=0$.
-The effects of different $\alpha$ on mixup degree $\lambda$ is discussed in Appendix.
+The effects of different $\alpha$ on mixup degree $\lambda$ is discussed in Appendix sec.
+4.3
 Figure 5: (a) Effect on learning rate on average test accuracy(out-of-sample) gap and on the cold-start users.
@@
 (Dynamic-only).
-In this scenario we used a fixed $Mu$ for all communication rounds.
+In this scenario we used a fixed MuMu for all communication rounds.
 (Sigmoid).
@@
 (Gradual).
-A simple linear function has been adapted for updating $Mu$.
+A simple linear function has been adapted for updating MuMu.
 (Random) Mixup degree $\lambda_{i}$ is selected randomly from $\beta$ distribution.
@@
 The results of this comparison are presented in Figure 5 (a).
-Additional details are provided in the Appendix 10.3.
+Additional details are provided in the Appendix sec.
+3.4.1.
+Analytical analysis of the Effect of learning rate and mixup degree.
+6.5 Communication and Computational Cost
+Although pMixFed introduces additional computation compared to partial personalization baselines (e.g., FedAlt, FedSim), we analyze its efficiency in terms of communication and runtime.
+Figure 6(a) shows the number of layers frozen at each communication round due to a zero mixup degree, which naturally reduces the amount of information exchanged.
+Figure 6(b) reports the average communication overhead across 100 clients for CNN and MobileNet on CIFAR-100.
+Figure 6: (a) Number of frozen layers per communication round due to zero mixup degree. (b) Average communication overhead per round across 100 clients for CNN and MobileNet on CIFAR-100.
+Compared to full-model FL (FedAvg), pMixFed does not increase the parameter size transmitted per layer.
+While its per-round communication cost is higher than partial methods (FedAlt, FedSim), two key properties compensate for this overhead:
+Faster convergence and early stopping.
+pMixFed reaches target accuracy in significantly fewer communication rounds, reducing the overall cost.
+In large-scale and highly heterogeneous settings (e.g., MobileNet on CIFAR-100), FedAlt and FedSim often diverge or require far more rounds to achieve comparable accuracy.
+Adaptive layer freezing.
+The mixup degree $\lambda_{i}^{(t)}$, updated dynamically via Eq.
+(5)–(6), drives progressive freezing of layers.
+As training advances, an increasing number of layers are excluded from both broadcasting and aggregation, further reducing transferred parameters in later rounds.
+We also evaluated computational efficiency by measuring GPU usage and runtime per round (Table 5).
+In the most demanding case (CIFAR-100, MobileNet, $C=100\%$), pMixFed is only slightly slower (1–7s per round) and requires marginally more GPU memory (1–22MB).
+Given its faster convergence (on average, $\sim$10% fewer rounds) and improved stability, the effective overhead is substantially mitigated.
+Table 5: GPU usage and computational time per global round using MobileNet on CIFAR-100 ($C=100\%$).
+Method $N=100$ $N=10$
+Time (s) Memory (MB) Time (s) Memory (MB)
+FedAvg 94.16 5362.44 95.31 735.25
+FedAlt 94.75 5372.77 92.86 746.85
+FedSim 95.72 5378.19 91.43 746.85
+pMixFed 100.03 5376.90 98.48 762.99
+In summary, while pMixFed incurs a minor per-round overhead, its faster convergence and adaptive freezing mechanism reduce the overall communication and computation cost, making it more scalable and robust in heterogeneous FL settings.
 7 Conclusions
@@
 Our approach features adaptive partitioning between shared and personalized layers, along with a gradual transition for personalization, enabling seamless adaptation for local clients, improved generalization across clients, and reduced risk of catastrophic forgetting.
+It should be noted that while Mixup has previously been used in FL (e.g., XORMixup, FEDMIX, FedMix), its role was limited to data augmentation.
+Similarly, model partitioning has been explored in works like FedAlt and FedSim, but only with fixed cut-off layers.
+In contrast, pMixFed introduces a novel, adaptive integration of layer-wise Mixup directly within the model parameter space—not the feature space.
+This integration is dynamically updated every communication round based on each client’s local accuracy (personalization) relative to the global accuracy (generalization).
+Governed by a tunable parameter $\mu$, pMixFed enables gradual, client-specific personalization across layers, mitigating global-local discrepancies and client drift in a fully dynamic manner.
+These innovations distinguish pMixFed from prior approaches, which rely on static strategies.
 We provided a theoretical analysis of pMixFed to study the properties of its convergence.
@@
 Empirically, pMixFed exhibited faster training times, increased robustness, and better handling of data heterogeneity compared to state-of-the-art PFL models.
-Future research directions include exploring multi-Modal personalization and adapting pMixFed for working on resource-constrained devices
-7 Acknowledgment
-This work is based upon the work partly supported by the National Center for Transportation Cybersecurity and Resiliency (TraCR) (a U.S.
+Future research directions include exploring multi-Modal personalization, exploiting other mixup variants within the parameter space and adapting pMixFed as a dynamic scheduling technique in resource-limited distributed setting.
+Acknowledgments and Disclosure of Funding
+This work was partly supported by the National Center for Transportation Cybersecurity and Resiliency (TraCR) (a U.S.
 Department of Transportation National University Transportation Center) headquartered at Clemson University, Clemson, South Carolina, USA.
@@
 8.1 Dynamic Mixup Degree
-Among the various partitioning strategies for partial PFL discussed in the main paper and in [34], one of the most widely adopted techniques is assigning higher layers of local model $L_{l,k}^{t}$ to personalization while allowing the base layers $L_{g,k}^{t}$ to be shared across clients as the global model [33, 45, 4].
-This design aligns with insights from the Model-Agnostic Meta-Learning (MAML) algorithm [10], which demonstrates that lower layers generally retain task-agnostic, generalized features, while the higher layers capture task-specific, personalized characteristics.
-Accordingly, in this work, we designate the head of the model as containing personalized information, while the base layers represent generalized information shared across clients  [^4] It should be noted that our method is fully adaptable to different partial PFL designs as well discussed in [34]..
-Broadcasting: To achieve a nuanced and gradual transition in the mixing process between the global model and the local model, we define the mix degree $\lambda_{i}$ as follows.
+Among the various partitioning strategies for partial PFL discussed in the main paper and in pillutla2022federated, one of the most widely adopted techniques is assigning higher layers of local model $L_{l,k}^{t}$ to personalization while allowing the base layers $L_{g,k}^{t}$ to be shared across clients as the global model oh2021fedbabu; sun2023fedperfix; arivazhagan2019federated.
+This design aligns with insights from the Model-Agnostic Meta-Learning (MAML) algorithm collins2022maml, which demonstrates that lower layers generally retain task-agnostic, generalized features, while the higher layers capture task-specific, personalized characteristics.
+Accordingly, in this work, we designate the head of the model as containing personalized information, while the base layers represent generalized information shared across clients  [^6] It should be noted that our method is fully adaptable to different partial PFL designs as well discussed in pillutla2022federated..
+8.1.1 Broadcasting:
+To achieve a nuanced and gradual transition in the mixing process between the global model and the local model, we define the mix degree $\lambda_{i}$ as follows.
 The local model’s head $L_{n}$ remains frozen, and the head of the global model is excluded from being shared with the local model.
@@
 The darker color shows the higher mixup degree ($\lambda_{i}$) for the corresponding layer $i$.
-Aggregation: The aggregation stage focuses on preserving the generalized information from the history of previous global models to mitigate the catastrophic forgetting problem.
+8.1.2 Aggregation:
+The aggregation stage focuses on preserving the generalized information from the history of previous global models to mitigate the catastrophic forgetting problem.
 In contrast to the broadcasting process, the primary goal here is to retain the generalized information of the previous global model, which encapsulates the history of all prior models $H\big|_{1}^{t-1}{G}$.
@@
 To enable online, adaptive updates of the mix degree $\lambda_{i}$, we implemented several algorithms, as detailed in Section 5.4 Ablation Study.
-We observed a clear relationship between the mixup degree and the similar behavior to lr, which is further discussed in Section 9.4.1.
+We observed a clear relationship between the mixup degree and the similar behavior to lr, which is further discussed in Section 4.
 Inspired by learning rate schedulers, we introduced the mix-factor $\mu$ to adaptively update the layer-wise mixup degree based on the current communication round $\delta=t/T$ and the relative performance of the current local model $L_{k}^{t}$ compared to the global model $G^{t}$.
@@
 Specifically, if the local model accuracy $Acc_{l}$ significantly exceeds that of the global model ($Acc_{G}$) in the current round ($t$) such that $Acc\gg 1$, less information is shared from the global model.
-Conversely, when $10<Acc<$ 1, the global model dominates the parameter updates in both the global and local models.
+Conversely, when $0<Acc<1$, the global model dominates the parameter updates in both the global and local models.
 Figure 8 shows the distribution of the calculated $\mu$ across different communication rounds for client 0 in the broadcasting stage.
@@
 For instance, if only 40% of clients have more than 4 layers, the generalization degree (in both broadcasting and aggregation stages), will be less than $0.4$ in each training round due to different participation rates.
+8.4 Difference from other works:
+While Mixup has previously been used in FL (e.g., XORMixup, FEDMIX, FedMix), its role was limited to data augmentation.
+Similarly, model partitioning has been explored in works like FedAlt and FedSim, but only with fixed cut-off layers.
+In contrast, pMixFed introduces a novel, adaptive integration of layer-wise Mixup directly within the model parameter space—not the feature space.
+This integration is dynamically updated every communication round based on each client’s local accuracy (personalization) relative to the global accuracy (generalization).
+Governed by a tunable parameter $\mu$, pMixFed enables gradual, client-specific personalization across layers, mitigating global-local discrepancies and client drift in a fully dynamic manner.
+These innovations distinguish pMixFed from prior approaches, which rely on static strategies.
 Figure 8: mix-factors for client 0 in different communication rounds
-9 Theoretical Analysis
-In this section, we provide the convergence analysis of pMixFed.
-Moreover, We compare the aggregation process and the server global model update of the FedSGD algorithm with our proposed mixed aggregation stage in pMixFed.
-We begin by introducing the key notations and assumptions used throughout the convergence analysis.
-Notations:
-  • $t\in\{0,\ldots,T-1\}$: communication round index.
-  • $\eta_{l}$: learning rate for local update, $\eta_{g}$: learning rate for global updates.
-  • $\lambda_{k,i}$: mixup coefficient for client $k$ at layer $i$ in round $t$, $\lambda_{k}$: mixup coefficient for client $k$ in round $t$ (assuming uniform across layers).
-  • $G^{(t)}$: global model parameters at round $t$, $L_{k}^{(t)}$: local model parameters of client $k$ at round $t$.
-  • $|\mathcal{D}_{k}|$: size of the dataset at client $k$, $|\mathcal{D}|$: total size of datasets across all clients.
-  • $\nabla L_{k}^{(t+1)}$: gradient of the local model at client $k$ in round $t$.
-We make the following assumptions to establish the convergence properties of pMixFed:
-Assumption 1.
-Unbiased Gradient Estimation with Bounded Variance. The gradient estimate for local model updates is unbiased and has a bounded variance, i.e.,
-$\mathbb{E}[\nabla\mathcal{L}_{k}^{(t)}]=\nabla\mathcal{L}_{k}^{(t)},\quad\text{Var}(\nabla\mathcal{L}_{k}^{(t)})\leq\sigma^{2}.$ (9)
-Assumption 2 (Smoothness of Local Objectives).
-The local objective functions $L_{k}(\cdot)$ are $L$-smooth, i.e.,
-$\small\|\nabla L_{k}^{(t+1)}-\nabla L_{k}^{(t)}\|\leq L\|L_{k}^{(t+1)}-L_{k}^{(t)}\|.$ (10)
-Assumption 3 (Bounded Gradients).
-The gradients at each client are bounded, i.e.,
-$\small\|\nabla L_{k}^{(t+1)}\|\leq G,\quad\forall k,t.$ (11)
-These assumptions are standard in convergence analysis and ensure that the optimization process is well-behaved.
-9.1 Local Training Convergence
-We first analyze the local training updates that occur between communication rounds.
-Lemma 1.
-Local Model Training Progress. Under Assumptions 2 and 1, after $r$ local updates, the expected loss for client $k$ satisfies:
-$\mathbb{E}[\mathcal{L}_{k}^{(t+r)}]\leq\mathcal{L}_{k}^{(t)}+\left(\frac{L_{1}\eta_{l}^{2}}{2}-\eta_{l}\right)\sum_{j=0}^{r-1}\|\nabla\mathcal{L}_{k}^{(t+j)}\|^{2}+\frac{L_{1}r\eta_{l}^{2}\sigma^{2}}{2}.$ (12)
-This lemma provides a bound on how the local training process improves the loss function.
-The bound depends on the learning rate $\eta_{l}$, the smoothness constant $L_{1}$, and the variance of the gradient estimates $\sigma^{2}$.
 Figure 9: Adaptive mix-factor $\mu_{k}^{t}$ according to the accuracy ration $Acc^{t}=Acc^{t}_{L}/Acc^{t}_{G}$ in different communication rounds.
-9.2 Global Model Aggregation
-Next, we consider the effect of aggregating local models at the server after each communication round.
-Lemma 2.
-Global Model Aggregation Dynamics. After aggregating the local models, the global model loss changes as follows:
-$\mathbb{E}[\mathcal{L}^{(t+1)}]\leq\mathbb{E}[\mathcal{L}^{(t)}]+\eta_{g}\delta^{2},$ (13)
-where $\delta^{2}$ bounds the difference between the global and local model parameters before and after aggregation.
-This lemma shows that, during aggregation, the global loss increases by a term proportional to the global learning rate $\eta_{g}$ and the parameter variation $\delta^{2}$.
-9.3 Convergence of pMixFed
-With the above lemmas, we can now derive the convergence properties of pMixFed.
-Theorem 1.
-Convergence Rate of pMixFed. Under the previously stated assumptions, the expected global loss after a complete round of communication and local training satisfies:
-$\small\mathbb{E}[\mathcal{L}^{(t+1)}]\leq\mathcal{L}^{(t)}+\left(\frac{L_{1}\eta_{l}^{2}}{2}-\eta_{l}\right)\sum_{k=1}^{K}\sum_{j=0}^{r-1}\|\nabla\mathcal{L}_{k}^{(t+j)}\|^{2}+\frac{L_{1}r\eta_{l}^{2}\sigma^{2}}{2}+\eta_{g}\delta^{2}.$ (14)
-This theorem demonstrates that the global loss decreases with each communication round, with the convergence depending on the local and global learning rates, the gradient norms, and the parameter variation.
-Theorem 2.
-Non-Convex Convergence Rate. For non-convex loss functions, pMixFed achieves convergence at the following rate:
-$\frac{1}{T}\sum_{t=0}^{T-1}\sum_{k=1}^{K}\|\nabla\mathcal{L}_{k}^{(t)}\|^{2}\leq\mathcal{O}(1/T),$ (15)
-where $T$ is the total number of communication rounds.
-This final theorem indicates that pMixFed achieves non-convex convergence with a rate of $\mathcal{O}(1/T)$, demonstrating that the algorithm improves over time as the number of communication rounds increases.
-9.4 The effect of Aggregating model parameters and gradients on catastrophic forgetting
-In FedSGD, the gradients are aggregated and the server will be update the global model according to the aggregated gradients.
-the FedSGD is sometimes preferred over FedAvg due to its potentially faster convergence.
-However, it lacks robustness in heterogeneous environments.
-pMixFed leverages the faster convergence characteristics of FedSGD by incorporating early stopping mechanisms, facilitated by the use of mixup.
-As demonstrated in Section 5, the mixup factor $\lambda$ functions analogously to an SGD update at the server, even though pMixFed aggregates model weights rather than gradients, similar to FedAvg.
-Section 9.4 provides a more detailed explanation of this mechanism.
-the pMixFed algorithm combines the advantages of FedSGD and FedAvg by aggregating model weights rather than gradients, while still ensuring convergence even in heterogeneous data settings.
-The incorporation of the mixup mechanism enhances stability, providing faster convergence rates compared to FedSGD, particularly in non-convex settings.
-Considering local SGD steps $r$ and $\eta_{l}$ as the local learning rate we can further expand the above formulation:
-$\begin{split}&G^{t+1}=G^{t}-\eta_{g}\sum_{i=1}^{k}\Omega_{k}\left[L_{k}^{(t+r)}-L_{k}^{(t)}\right]\\ &G^{t+1}=G^{t}-\eta_{g}.\eta_{l}\sum_{i=1}^{k}\Omega_{k}\sum_{j=0}^{r-1}\nabla{L_{k}^{(t+j)}}\end{split}$ (16)
-Since in FedSGD, the global model $G^{(t)}$ is fully shared with each local model $L_{k}^{(t)}$ in each communication round t, and $\sum_{i=1}^{k}\Omega_{k}=1$, we can update the problem as a constrained optimization problem.
-$G^{t+1}=G^{t}-\eta_{g}.\eta_{l}\ \sum_{j=1}^{r-1}\nabla{G^{(t+j)}}$
-$\sum_{j=2}^{r-1}\nabla{G^{(t+j)}=(1-\eta_{g}.\eta_{l})G^{t}-(1+\eta_{g}.\eta_{l})G^{t+1}\ }$
-The right term of the equation is the history of global model gradients which could be written as $H\big{|}_{2}^{r-1}\nabla{G^{t}}=a.G^{t}-b.G^{t+1}$.
-The problem of the FedSGD algorithm is that the gradients are too small to keep the state of the previous global model, leading to catastrophic forgetting in partial models that use gradients in their aggregation stage.
-We hypothesize that the mixup coefficient $\lambda$ acts similarly to the learning rate $\eta$, suggesting that $\lambda$ plays a role analogous to $\eta$ in FedSGD.
-FedSGD Update Rule : In FedSGD, the global model $G^{(t)}$ is updated by aggregating the gradients from all clients:
-$\small G^{(t+1)}=G^{(t)}-\eta_{g}\sum_{k=1}^{K}\Omega_{k}\nabla L_{k}^{(t+1)},$ (17)
-where $\Omega_{k}=\frac{|\mathcal{D}_{k}|}{|\mathcal{D}|}$ is the weight associated with client $k$. Since $L_{k}^{(t+1)}=L_{k}^{(t)}-\eta_{l}\nabla L_{k}^{(t+1)}$, we can rewrite the update as:
-$\small G^{(t+1)}=G^{(t)}-\eta_{g}\sum_{k=1}^{K}\Omega_{k}\left(\frac{L_{k}^{(t)}-L_{k}^{(t+1)}}{\eta_{l}}\right).$ (18)
-pMixFed Update Rule: In pMixFed, the global model update incorporates mixup coefficients:
-$\small G^{(t+1)}=\sum_{k=1}^{K}\Omega_{k}\left[(1-\lambda_{k})L_{k}^{(t+1)}+\lambda_{k}G^{(t)}\right].$ (19)
-Assuming $\lambda_{k,i}=\lambda_{k}$ for all layers $i$, and considering that $G^{(t)}$ is sent to all clients at round $t$, we simplify the update to:
-$\small G^{(t+1)}=(1-\lambda_{k})\sum_{k=1}^{K}\Omega_{k}L_{k}^{(t+1)}+\lambda_{k}G^{(t)}.$ (20)
-9.4.1 Analytical analysis of the Effect of learning rate and mixup degree
-To establish the relationship between $\lambda$ and $\eta_{g}$, we align the FedSGD and pMixFed update equations. From Equation (20), rearranged:
-$\small G^{(t+1)}-\lambda_{k}G^{(t)}=(1-\lambda_{k})\sum_{k=1}^{K}\Omega_{k}L_{k}^{(t+1)}.$ (21)
-From Equation (18), rearranged:
-$\small G^{(t+1)}=G^{(t)}-\frac{\eta_{g}}{\eta_{l}}\sum_{k=1}^{K}\Omega_{k}\left(L_{k}^{(t)}-L_{k}^{(t+1)}\right).$ (22)
-Assuming $L_{k}^{(t)}$ is replaced with $G^{(t)}$ at round $t$ for FedSGD, we have $L_{k}^{(t)}=G^{(t)}$. Substituting this into Equation (22):
-$G^{(t+1)}=G^{(t)}-\frac{\eta_{g}}{\eta_{l}}\sum_{k=1}^{K}\Omega_{k}\left(G^{(t)}-L_{k}^{(t+1)}\right).$ (23)
-Simplifying and Comparing with Equation (20), we see that if: $\lambda_{k}=\frac{\eta_{g}}{\eta_{l}}$. then the updates are analogous.
-$G^{(t+1)}=G^{(t)}\left(1-\frac{\eta_{g}}{\eta_{l}}\sum_{k=1}^{K}\Omega_{k}\right)+\frac{\eta_{g}}{\eta_{l}}\sum_{k=1}^{K}\Omega_{k}L_{k}^{(t+1)}.$ (24)
-Theorem 3.
-Under Assumptions 2 and 3, the mixup coefficient $\lambda_{k}$ in pMixFed acts similarly to the learning rate ratio $\frac{\eta_{g}}{\eta_{l}}$ in FedSGD, such that:$\lambda_{k}=\frac{\eta_{g}}{\eta_{l}}$.
-This implies that the mixup mechanism in pMixFed can be interpreted as a form of learning rate control analogous to FedSGD.
-Proof.
-Starting from the pMixFed update in Equation (21):
-$G^{(t+1)}-\lambda_{k}G^{(t)}=(1-\lambda_{k})\sum_{k=1}^{K}\Omega_{k}L_{k}^{(t+1)}.$ (25)
-From the rearranged FedSGD update in Equation (24):
-$G^{(t+1)}=G^{(t)}\left(1-\frac{\eta_{g}}{\eta_{l}}\right)+\frac{\eta_{g}}{\eta_{l}}\sum_{k=1}^{K}\Omega_{k}L_{k}^{(t+1)}.$ (26)
-Setting the two expressions equal:
-$G^{(t)}\left(1-\lambda_{k}\right)+(1-\lambda_{k})\sum_{k=1}^{K}\Omega_{k}L_{k}^{(t+1)}=$
-$G^{(t)}\left(1-\frac{\eta_{g}}{\eta_{l}}\right)+\frac{\eta_{g}}{\eta_{l}}\sum_{k=1}^{K}\Omega_{k}L_{k}^{(t+1)}.$ (27)
-This simplifies to:
-$\lambda_{k}=\frac{\eta_{g}}{\eta_{l}}.$ (28)
-Given that $\lambda_{k}$ must lie in the range $[0,1]$, this relationship holds when $\eta_{g}\leq\eta_{l}$.
-Since both $\eta_{g}$ and $\eta_{l}$ are also constrained within the interval $[0,1]$, our findings are consistent with previous studies [33, 10], which recommend keeping the head unfrozen (i.e., $\eta_{l}\neq 0$).
-Remark1.
-Theorem 3 establishes a direct relationship between the mixup coefficient $\lambda_{k}$ and the learning rates used in local and global updates.
-This insight allows us to interpret the mixup mechanism in pMixFed as adjusting the effective learning rate at the server, providing a theoretical foundation for selecting $\lambda_{k}$ based on desired convergence properties.
-Remark2.
-practice, this relationship suggests that by tuning $\lambda_{k}$, we can control the influence of the global model versus the local models in the aggregation process, similar to adjusting the learning rate in FedSGD.
-This is particularly beneficial in heterogeneous environments where clients may have varying data distributions.
-Our theoretical analysis indicates that the mixup coefficient $\lambda_{k}$ in pMixFed plays a role analogous to the learning rate in FedSGD.
-This equivalence provides a deeper understanding of how pMixFed leverages the strengths of FedSGD while mitigating its weaknesses in heterogeneous settings.
-By appropriately choosing $\lambda_{k}$, pMixFed can achieve faster convergence and improved robustness.
+9 Proof of Theorems:
+Proof of Theorem 1 (SGD form with an effective step).
+By definition of the pMixFed aggregation,
+$\theta^{t+1}\penalty 10000\ =\penalty 10000\ \sum_{k\in\mathcal{U}_{t}}\Omega_{k}\Big(\lambda_{k}^{t}\,\theta^{t}+(1-\lambda_{k}^{t})\,\theta_{k}^{t+1}\Big)\penalty 10000\ =\penalty 10000\ \bar{\lambda}^{\,t}\,\theta^{t}+(1-\bar{\lambda}^{\,t})\sum_{k\in\mathcal{U}_{t}}\Omega_{k}\theta_{k}^{t+1},$
+where $\bar{\lambda}^{\,t}:=\sum_{k\in\mathcal{U}_{t}}\Omega_{k}\lambda_{k}^{t}\in[0,1]$.
+Each selected client performs one local SGD step from the broadcasted model $\theta^{t}$: $\theta_{k}^{t+1}=\theta^{t}-\eta_{\ell}\,g_{k}^{t}$, where $g_{k}^{t}$ is the stochastic gradient computed on client $k$ at $\theta^{t}$.
+Substituting into the aggregation gives
+$\theta^{t+1}\penalty 10000\ =\penalty 10000\ \bar{\lambda}^{\,t}\,\theta^{t}\penalty 10000\ +\penalty 10000\ (1-\bar{\lambda}^{\,t})\sum_{k\in\mathcal{U}_{t}}\Omega_{k}\big(\theta^{t}-\eta_{\ell}g_{k}^{t}\big)\penalty 10000\ =\penalty 10000\ \theta^{t}-(1-\bar{\lambda}^{\,t})\,\eta_{\ell}\sum_{k\in\mathcal{U}_{t}}\Omega_{k}\,g_{k}^{t}.$
+Therefore the server update is exactly an SGD step on $F(\theta)=\sum_{k}\Omega_{k}F_{k}(\theta)$ with the effective step size $\eta_{\mathrm{eff}}^{\,t}:=(1-\bar{\lambda}^{\,t})\,\eta_{\ell}$, i.e., $\theta^{t+1}=\theta^{t}-\eta_{\mathrm{eff}}^{\,t}\sum_{k\in\mathcal{U}_{t}}\Omega_{k}\,g_{k}^{t}$.
+Proof of Theorem 2 (Coefficient matching with FedSGD).
+Assume the server writes its update as a FedSGD step with server stepsize $\eta_{g}>0$: $\theta^{t+1}=\theta^{t}-\eta_{g}\sum_{k\in\mathcal{U}_{t}}\Omega_{k}\,g_{k}^{t}$.
+From Theorem 1 we also have the pMixFed SGD form $\theta^{t+1}=\theta^{t}-(1-\bar{\lambda}^{\,t})\,\eta_{\ell}\sum_{k\in\mathcal{U}_{t}}\Omega_{k}\,g_{k}^{t}$.
+Equality of the two updates for arbitrary realizations of the stochastic gradients holds if and only if their coefficients on the aggregated gradient coincide, i.e.,
+[…]
+Since by construction $\bar{\lambda}^{\,t}\in[0,1]$, we immediately obtain the admissible range $0\leq\eta_{g}/\eta_{\ell}\leq 1$.
+This proves the necessary and sufficient condition for coefficient matching.
+Proof of Theorem 3 (Nonconvex rate).
+Recall the pMixFed update in SGD form
+$\theta^{t+1}\;=\;\theta^{t}\;-\;\eta_{\mathrm{eff}}\,\underbrace{\sum_{k\in\mathcal{U}_{t}}\Omega_{k}\,g_{k}^{t}}_{:=\,\bar{g}^{t}},\qquad\eta_{\mathrm{eff}}\;\leq\;\frac{1}{L},$
+where by (A2) the aggregated stochastic gradient $\bar{g}^{t}$ is unbiased, $\mathbb{E}[\bar{g}^{t}\mid\theta^{t}]=\nabla F(\theta^{t})$, and has bounded second moment $\mathbb{E}\|\bar{g}^{t}-\nabla F(\theta^{t})\|^{2}\leq\sigma^{2}$.
+By $L$-smoothness (A1),
+$F(\theta^{t+1})\;\leq\;F(\theta^{t})+\big\langle\nabla F(\theta^{t}),\,\theta^{t+1}-\theta^{t}\big\rangle+\frac{L}{2}\,\|\theta^{t+1}-\theta^{t}\|^{2}.$
+Substitute $\theta^{t+1}-\theta^{t}=-\eta_{\mathrm{eff}}\,\bar{g}^{t}$, take conditional expectation given $\theta^{t}$, and use $\mathbb{E}\|\bar{g}^{t}\|^{2}=\|\nabla F(\theta^{t})\|^{2}+\mathbb{E}\|\bar{g}^{t}-\nabla F(\theta^{t})\|^{2}\leq\|\nabla F(\theta^{t})\|^{2}+\sigma^{2}$ to obtain
+$\mathbb{E}\!\left[F(\theta^{t+1})\,\middle|\,\theta^{t}\right]\;\leq\;F(\theta^{t})\;-\;\eta_{\mathrm{eff}}\,\|\nabla F(\theta^{t})\|^{2}\;+\;\frac{L\,\eta_{\mathrm{eff}}^{2}}{2}\Big(\|\nabla F(\theta^{t})\|^{2}+\sigma^{2}\Big).$
+Rearranging,
+$\mathbb{E}\!\left[F(\theta^{t+1})\,\middle|\,\theta^{t}\right]\;\leq\;F(\theta^{t})\;-\;\Big(\eta_{\mathrm{eff}}-\tfrac{L\eta_{\mathrm{eff}}^{2}}{2}\Big)\,\|\nabla F(\theta^{t})\|^{2}\;+\;\frac{L\,\eta_{\mathrm{eff}}^{2}}{2}\,\sigma^{2}.$
+Since $\eta_{\mathrm{eff}}\leq 1/L$, we have $\eta_{\mathrm{eff}}-\tfrac{L\eta_{\mathrm{eff}}^{2}}{2}\geq\tfrac{\eta_{\mathrm{eff}}}{2}$; taking total expectation yields the one-step descent inequality
+$\mathbb{E}\big[F(\theta^{t+1})\big]\;\leq\;\mathbb{E}\big[F(\theta^{t})\big]\;-\;\frac{\eta_{\mathrm{eff}}}{2}\,\mathbb{E}\big\|\nabla F(\theta^{t})\big\|^{2}\;+\;\frac{L\,\eta_{\mathrm{eff}}^{2}}{2}\,\sigma^{2}.$
+Summing $(\star)$ over $t=0,\dots,T-1$ and telescoping gives
+$\frac{\eta_{\mathrm{eff}}}{2}\sum_{t=0}^{T-1}\mathbb{E}\big\|\nabla F(\theta^{t})\big\|^{2}\;\leq\;F(\theta^{0})-F^{\star}\;+\;\frac{L\,\eta_{\mathrm{eff}}^{2}}{2}\,T\,\sigma^{2}.$
+Divide both sides by $T\,\eta_{\mathrm{eff}}/2$ to conclude
+$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\big\|\nabla F(\theta^{t})\big\|^{2}\;\leq\;\frac{2\,(F(\theta^{0})-F^{\star})}{T\,\eta_{\mathrm{eff}}}\;+\;L\,\eta_{\mathrm{eff}}\,\sigma^{2},$
+which is precisely the claimed nonconvex rate.
+Proof of Theorem 4 (Strongly convex case).
+Assume (A1)–(A3) and $\eta_{\mathrm{eff}}\leq\min\{1/L,\,1/(2\mu)\}$. Starting again from the smoothness inequality and the SGD update,
+$\mathbb{E}\big[F(\theta^{t+1})\,\big|\,\theta^{t}\big]\;\leq\;F(\theta^{t})\;-\;\eta_{\mathrm{eff}}\,\|\nabla F(\theta^{t})\|^{2}\;+\;\frac{L\,\eta_{\mathrm{eff}}^{2}}{2}\Big(\|\nabla F(\theta^{t})\|^{2}+\sigma^{2}\Big).$
+As before, with $\eta_{\mathrm{eff}}\leq 1/L$ we get
+$\mathbb{E}\big[F(\theta^{t+1})\,\big|\,\theta^{t}\big]\;\leq\;F(\theta^{t})\;-\;\frac{\eta_{\mathrm{eff}}}{2}\,\|\nabla F(\theta^{t})\|^{2}\;+\;\frac{L\,\eta_{\mathrm{eff}}^{2}}{2}\,\sigma^{2}.$
+Strong convexity implies the Polyak–Łojasiewicz (PL) inequality $\|\nabla F(\theta^{t})\|^{2}\geq 2\mu\big(F(\theta^{t})-F^{\star}\big)$.
+Plugging this into $(\dagger)$ and taking total expectation gives
+$\mathbb{E}\big[F(\theta^{t+1})-F^{\star}\big]\;\leq\;\big(1-\mu\,\eta_{\mathrm{eff}}\big)\,\mathbb{E}\big[F(\theta^{t})-F^{\star}\big]\;+\;\frac{L\,\eta_{\mathrm{eff}}^{2}}{2}\,\sigma^{2}.$
+Unrolling the linear recursion,
+$\mathbb{E}\big[F(\theta^{t})-F^{\star}\big]\;\leq\;(1-\mu\,\eta_{\mathrm{eff}})^{t}\big(F(\theta^{0})-F^{\star}\big)\;+\;\frac{L\,\eta_{\mathrm{eff}}\,\sigma^{2}}{2\mu}.$
+Finally, strong convexity also yields $F(\theta)-F^{\star}\geq\tfrac{\mu}{2}\|\theta-\theta^{\star}\|^{2}$, so
+$\mathbb{E}\,\|\theta^{t}-\theta^{\star}\|^{2}\;\leq\;\frac{2}{\mu}\,\mathbb{E}\big[F(\theta^{t})-F^{\star}\big]\;\leq\;(1-\mu\,\eta_{\mathrm{eff}})^{t}\,\|\theta^{0}-\theta^{\star}\|^{2}\;+\;\mathcal{O}\!\Big(\tfrac{\sigma^{2}}{\mu}\,\eta_{\mathrm{eff}}\Big),$
+which is the claimed result.
 10 Additional Details about the Experiments
 10.1 Experimental Setup
-Dataset: We used three widely used federated learning datasets: MNIST [23], CIFAR-10, CIFAR-100 [3], and EMNIST [8].
-CIFAR-10 consists of 50,000 images of size $32\times 32$ for training and 10,000 images for testing.
-CIFAR-100, on the other hand, consists of 100 classes, with 500 $32\times 32$ images per class for training and 100 images per class for testing.
-MNIST contains 10 labels and includes 60,000 samples of $28\times 28$ grayscale images for training and 10,000 for testing.
-For creating heterogenity, we followed the
-Training Details: For evaluation, We have reported the average test accuracy of the global model [59] for different approaches.
+For creating heterogenity, we followed the Training Details: For evaluation, We have reported the average test accuracy of the global model yuan2021we for different approaches.
 The final global model at the last communication round is saved and used during the evaluation.
@@
 [^7] As discussed in the main paper, the change of hyper-parameters such as lr, batch size, momentum and even changing the optimizer to adam didn’t help with the performance drop in most cases.
+10.2 Caltech-101 Experimental Setup and Results
+Dataset.
+Caltech-101 fei2004learning contains 9,146 images across 101 object categories plus a background class, with $\sim$40–800 images per class (most near 50).
+Images are resized to $224\times 224$ for CNN training.
+We followed prior FL literature fallah2020personalized; yang2023fedrep; pillutla2022federated in simulating heterogeneous splits: each client receives at most $S=30$ classes sampled by Dirichlet distribution ($\alpha=0.5$).
+We evaluate four federated settings: $N=\{10,100\}$ clients with participation $C=\{10\%,100\%\}$.
+Model and Training.
+For consistency with CIFAR/MNIST, we adopt a CNN backbone (4 convolutional layers + 1 fully connected layer).
+Each client trains for $r=4$ local epochs per round with batch size 32.
+The total communication rounds are $T=50$, using Adam optimizer with learning rate $1\times 10^{-3}$.
+For partial baselines (FedAlt), the split layer is fixed at the midpoint.
+For Per-FedAvg, meta-initialization is updated with inner learning rate $0.01$ and outer learning rate $0.001$.
+FedRep uses a shared representation with personalized heads (dimension 256).
+Evaluation is done by fine-tuning on local test splits for 4 epochs.
+Results.
+Table 6 reports detailed top-1 accuracy (%) across four evaluation angles (A–D), along with the average.
+Across all federated settings, pMixFed achieves the best performance, with particularly strong improvements in the heterogeneous regime ($N=100$, $C=10\%$).
+FedRep performs second-best overall, confirming the utility of representation sharing, while Per-FedAvg is strong when client counts are small.
+FedAlt trails consistently, highlighting the limitations of fixed partitioning.
+10.3 Results on Caltech-101 dataset
+We additionally evaluate on Caltech-101 fei2004learning, following the same heterogeneous FL setup as CIFAR/MNIST.
+Each client receives at most $S=30$ classes, and data is split using a Dirichlet distribution ($\alpha=0.5$).
+We report accuracy across four evaluation angles (A–D) and their mean (AVG).
+Table 6: Performance (%) on Caltech-101 (CNN backbone). Rows A–D denote evaluation angles; AVG is the mean across A–D.
+$N=100,\penalty 10000\ C=100\%$
+Method A B C D AVG
+pMixFed (Ours) 78.4 80.2 79.1 79.5 79.3
+FedRep 75.8 78.5 77.2 76.6 77.0
+Per-FedAvg 75.1 77.2 76.0 77.5 76.5
+FedAlt 70.8 74.1 73.5 73.9 73.1
+$N=10,\penalty 10000\ C=100\%$
+Method A B C D AVG
+pMixFed (Ours) 80.9 83.8 82.9 81.7 82.4
+FedRep 78.6 80.9 80.4 80.5 80.1
+Per-FedAvg 79.0 81.8 82.1 81.0 81.0
+FedAlt 74.0 76.8 75.2 76.0 75.5
+$N=100,\penalty 10000\ C=10\%$
+Method A B C D AVG
+pMixFed (Ours) 74.6 77.8 76.9 74.9 76.1
+FedRep 72.9 74.6 74.5 74.8 74.2
+Per-FedAvg 72.1 74.0 73.8 74.1 73.5
+FedAlt 67.4 70.8 69.9 69.8 69.5
+$N=10,\penalty 10000\ C=10\%$
+Method A B C D AVG
+pMixFed (Ours) 77.2 80.1 79.8 79.0 79.0
+FedRep 75.1 77.3 77.2 77.8 76.9
+Per-FedAvg 74.2 76.5 76.0 77.9 76.2
+FedAlt 70.5 72.6 72.9 72.8 72.2
 10.4 Out-of-Sample and Participation Gap
@@
 The test accuracy of pFedMix, while affected under a 10% participation rate, benefits significantly from seeing more clients.
-Increased client participation directly improves accuracy, as observed in previous studies [34].
+Increased client participation directly improves accuracy, as observed in previous studies pillutla2022federated.
 10.5 Ablation Study : The Effects of Different alpha on Mixup Degree
@@
 The value of $\lambda$, used in Eq.
-4, is naturally sampled from this distribution.
+9, is naturally sampled from this distribution.
 By varying the parameter $\alpha$, we can adjust how much mixing occurs between the global model $G$ and the local model $L_{k}$.
@@
 During these stages, the parameter $\alpha$ is adaptively changed as follows:
-    $\alpha=\begin{cases}0.1,&\text{initial stage (early training)}\\ 100,&\text{middle stage (convergence phase)}\\ 10,&\text{final stage (fine-tuning)}\end{cases}$ (29)
+[…]
     This adaptive strategy mimics the behavior of the original pFedMix algorithm while also allowing for more controlled exploration of different mixup combinations.
@@
 The results shows that the adaptive sampling which creats a form scheduling for mixup degree shows promising result even compared to the original algorithm using adaptive $\mu$.
-Table 3: Accuracy of $pMixFed$ with Different $\alpha$ Values in the Beta Distribution
+Table 7: Accuracy of pMixFedpMixFed with Different $\alpha$ Values in the Beta Distribution
 Dataset $\alpha=0.1$ $\alpha=0.5$ $\alpha=1$ $\alpha=2$ $\alpha=5$
@@
 CIFAR-100 42.3 45.6 47.2 49.8 48.1
-Table 4: Accuracy of $pMixFed$ with Different $\alpha$ Values and Sampling Strategies
+Table 8: Accuracy of pMixFedpMixFed with Different $\alpha$ Values and Sampling Strategies
 Dataset Sampling Strategy Accuracy(%)